### v2

# Two-stage Estimation for Causal Inference Involving a Semi-continuous Exposure

Abstract
Methods for causal inference are well developed for binary and continuous exposures, but in many settings, the exposure has a substantial mass at zero—such exposures are called semi-continuous. We propose a general causal framework for such semi-continuous exposures, together with a novel two-stage estimation strategy. A two-part propensity structure is introduced for the semi-continuous exposure, with one component for exposure status (exposed vs unexposed) and another for the exposure level among those exposed, and incorporates both into a marginal structural model that disentangles the effects of exposure status and dose. The two-stage procedure sequentially targets the causal dose–response among exposed individuals and the causal effect of exposure status at a reference dose, allowing flexibility in the choice of propensity score methods in the second stage. We establish consistency and asymptotic normality for the resulting estimators, and characterise their limiting values under misspecification of the propensity score models. Simulation studies evaluate finite sample performance and robustness, and an application to a study of prenatal alcohol exposure and child cognition demonstrates how the proposed methods can be used to address a range of scientific questions about both exposure status and exposure intensity.
Keywords Augmented inverse probability weighting (AIPW) $\cdot$ Causal inference $\cdot$ Inverse probability weighting (IPW) $\cdot$ Propensity score (PS) regression adjustment $\cdot$ Semi-continuous exposure $\cdot$ Two-stage approach
1 Introduction
In many areas of public health research, the goal is to estimate causal effects of exposures on outcomes using observational data subject to confounding. Propensity scores summarise measured confounders for a binary exposure and can be used to mitigate bias in estimation of causal effects (Rubin, 1974; Rosenbaum and Rubin, 1983, 1984). For continuous exposures, the generalised propensity score is defined as the conditional density of the exposure given covariates, evaluated at the observed exposure (Hirano and Imbens, 2004; Imai and Van Dyk, 2004). Inverse density weighting can be used to estimate dose–response relationships, but the resulting estimators may be unstable when outliers are present and weights are highly variable (Naimi et al., 2014), and correctly modelling the exposure distribution can be challenging.
In many applications, however, the exposure is neither purely binary nor fully continuous. Samples often consist of both unexposed individuals and exposed individuals receiving varying doses, yielding a semi-continuous exposure with a point mass at zero and a continuous component among the exposed. Such structures arise, for example, in environmental health studies where pollutants may be absent at some locations but vary in concentration where present (Begu et al., 2016), and in studies of behaviours such as smoking or alcohol consumption where many individuals abstain but usage among those who do not can vary considerably. Yet despite their prevalence, semi-continuous exposures have received relatively little attention in causal inference using propensity score methods.
As a motivating example, we consider the challenge of assessing the effect of prenatal alcohol exposure (PAE) on child cognition. We analyse data from a Detroit longitudinal cohort of African-American children in which maternal alcohol use during pregnancy was measured as average daily ounces of absolute alcohol, yielding a semi-continuous exposure with a point mass at zero among non-drinkers and a continuous right-skewed distribution among drinkers (Jacobson et al., 2002). Although epidemiological studies have reported associations between PAE and children’s cognitive outcomes (Jacobson et al., 2004, 2011; Lewis et al., 2015, 2016), the precise nature of the dose–response effect remains unclear. Scientific and public health inquiries focus on the potential adverse effects of any alcohol consumption during pregnancy, as well as how those risks escalate with increased consumption. To address such semi-continuous exposures, Akkaya Hocagil et al. (2021) construct a two-part propensity score, modelling exposure status via logistic regression and the dose component among the exposed via log linear regression. In related work, Li et al. (2023) outline a two-stage analysis for semi-continuous exposures with continuous outcomes, focusing on the modelling strategy and empirical illustration rather than on a formal counterfactual causal framework or asymptotic and robustness properties. We return to this study in Section 6.
In this article, we develop and evaluate a two-stage procedure for assessing the causal effect of a semi-continuous exposure on a continuous outcome. In Stage I, we use propensity score regression adjustment among the exposed to estimate the effect of the continuous exposure level. In Stage II, we assess the effect of exposure at a specified reference dose compared with no exposure, using a second propensity score defined for the binary exposure indicator. This two-stage structure enables a doubly robust procedure in Stage II through augmented inverse probability weighted (AIPW) estimating equations (Bang and Robins, 2005; Robins et al., 2007). We allow three distinct sets of confounders: one affecting exposure status (exposed versus unexposed), a second affecting the exposure dose among those exposed, and a third affecting both exposure status and dose. Within the potential outcomes framework, we define causal estimands for both status and dose, specify the assumptions required for identification, and derive the large sample properties of the resulting estimators. Extensive simulation studies are used to assess finite sample performance and to examine the impact of various forms of propensity score misspecification. Finally, we illustrate the methods using data from a study of PAE and child cognition.
The remainder of this article is organised as follows. In Section 2, we introduce notation, models, and assumptions for causal analysis with a semi-continuous exposure, and describe regression adjustment based on two propensity scores for the causal estimands of interest. In Section 3, we present the two-stage estimators for the binary status and continuous dose effects, and derive their asymptotic properties. In Section 4, we study the impact of propensity score misspecification by evaluating the limiting values of the estimators. Simulation studies are reported in Section 5, and in Section 6 we apply the methods to data from a study by Jacobson et al. (2002) on the effect of PAE on child cognition. Conclusions and future directions are discussed in Section 7.
2 Causal Analysis in a Two-part Model
2.1 Notation and Assumptions
Let $Y$ denote a continuous response variable, $T$ denote a non-negative random variable which represents the exposure (e.g. the ounces of absolute alcohol per day) of an individual, and let $A=I(T>0)$ indicate the exposure status. A $k_{1}\times 1$ vector of confounding variables $\mathbf{X}_{1}=(X_{11},...,X_{1k_{1}})^{\prime}$ has effects on $T$ and $Y$, a $k_{2}\times 1$ vector of $\mathbf{X}_{2}=(X_{21},...,X_{2k_{2}})^{\prime}$ confounds the relationship between the binary indicator $A$ and $Y$ among those with $A=1$, and a $k_{3}\times 1$ vector of confounders $\mathbf{X}_{3}=(X_{31},...,X_{3k_{3}})^{\prime}$ affects on $(A,T)$, and $Y$; see the directed acyclic graph in Figure 1. We then let $\mathbf{X}=(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime},\mathbf{X}_{3}^{\prime})^{\prime}$ be the $k\times 1$ full vector of confounders with $k=k_{1}+k_{2}+k_{3}$, $\mathbf{Z}_{1}=(\mathbf{X}_{1}^{\prime},\mathbf{X}_{3}^{\prime})^{\prime}$ be a $(k_{1}+k_{3})\times 1$ vector of confounders associated with the continuous exposure $T$ and $Y$ given $A=1$, and $\mathbf{Z}_{2}=(\mathbf{X}_{2}^{\prime},\mathbf{X}_{3}^{\prime})^{\prime}$ be a $(k_{2}+k_{3})\times 1$ vector of confounders associated with the binary exposure $A$ and the response. For a sample of $n$ independent individuals, the observed data is $\{(Y_{i},T_{i},A_{i},\mathbf{X}_{i}^{\prime}),i=1,\ldots,n\}$.
[…]
Figure 1: A directed acyclic graph for a two-part exposure model
As is common in dose-response modelling, we consider a log transformation of the continuous positive exposure $T$ to define the dose $D=\log T$, with $d$ its realized value. We then let $Y_{i}(1,d)$ represent the potential outcome for the $i$-th individual exposed at dose $d$ (Rubin, 1974; Splawa-Neyman et al., 1990); if the $i$-th individual is unexposed, then the dose is undefined, but we denote the potential outcome as $Y_{i}(0,d)$ for consistency of notation, taking it as understood that $d$ is undefined. For a semi-continuous exposure among those exposed, the average causal effect of a one-unit increase in dose from $d$ to $d+1$ is
$E\left\{Y(1,d+1)\right\}-E\left\{Y(1,d)\right\};$
we refer to this as the causal dose-response effect. To define the average causal effect of the binary exposure status we must specify a reference dose $c$ for the exposed individuals. We then let
$E\left\{Y(1,c)\right\}-E\left\{Y(0,d)\right\}$
denote the average causal effect of exposure at level $c$ corresponds to those that are unexposed. In a sample, we often use the sample average of the log volume among the exposed group for $c$, but any reference value can of course be set.
To identify the causal effects we extend the standard causal inference assumptions (Rubin, 1980, 1990; Rosenbaum and Rubin, 1983; Cole and Hernán, 2008) to the setting of a semi-continuous exposure; Appendix A provides a detailed discussion.
Assumption 1 (Stable Unit Treatment Value Assumption)
This assumption states that the potential outcome for one individual is independent of the exposure of another individual.
Assumption 2 (Consistency)
$Y_{i}(1,d)=Y_{i}$ if the $i$-th individual is exposed at dose $D_{i}=d$ and $Y_{i}(0,d)=Y_{i}$ if unexposed for all $d$ and $i=1,\ldots,n$.
Assumption 3 (Ignorability)
$Y_{i}(1,d)\perp D_{i}|A_{i}=1,\mathbf{X}_{i}=\mathbf{x}$ and $Y_{i}(1,d),Y_{i}(0,0)\perp A_{i}|\mathbf{X}_{i}$ for all $d$ and $i=1,\ldots,n$.
Assumption 4 (Positivity)
$f(D_{i}=d|A_{i}=1,\mathbf{Z}_{i1})>0$ for all $d$, and $P(A_{i}=a|\mathbf{Z}_{i2})>0$ for $a=0,1$ and $i=1,\ldots,n$.
2.2 Model Settings
Suppressing the index $i$ for individuals, we assume the data generation model for $Y$:
$Y=\theta_{0}+{\theta}_{11}A+{\theta}_{12}A(D-c)+\mathbf{X}^{\prime}_{1}\boldsymbol{\theta}_{21}+\mathbf{X}^{\prime}_{2}\boldsymbol{\theta}_{22}+\mathbf{X}^{\prime}_{3}\boldsymbol{\theta}_{23}+E,$ (1)
with the error $E\sim N(0,\tau^{2})$ and $E\perp(A,AD,\mathbf{X})$, $c$ a specified reference dose, $\boldsymbol{\theta}=(\theta_{0},{\theta}_{11},{\theta}_{12},\boldsymbol{\theta}_{2}^{\prime})^{\prime}$ is a $(k+3)\times 1$ vector of parameters with $\boldsymbol{\theta}_{21}=(\theta_{211},...,\theta_{21k_{1}})^{\prime}$, $\boldsymbol{\theta}_{22}=(\theta_{221},...,\theta_{22k_{2}})^{\prime}$, $\boldsymbol{\theta}_{23}=(\theta_{231},...,\theta_{23k_{3}})^{\prime}$, and $\boldsymbol{\theta}_{2}=(\boldsymbol{\theta}^{\prime}_{21},\boldsymbol{\theta}^{\prime}_{22},\boldsymbol{\theta}^{\prime}_{23})^{\prime}$; see the DAG of Figure 1. Note $\theta_{11}$ and $\theta_{12}$ in (1) represent the causal effects of the exposure at the reference dose (comparing to those who are unexposed) and the causal dose-response effect respectively. Since the mean response varies as a function of dose, a reference dose $c$ is required to define the parameter $\theta_{11}$; any value can be specified but values near the center of the distribution of doses among those exposed will often be appealing.
Here we consider a two-part model for the semi-continuous exposure with one for the exposure status and one for the dose (Smith et al., 2017). The exposure model for the dose among those exposed has the form
$E(D|A=1,\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})=\bar{\mathbf{Z}}^{\prime}_{1}\boldsymbol{\alpha}_{1}$ (2)
where $\bar{\mathbf{Z}}_{1}=(1,\mathbf{Z}_{1}^{\prime})^{\prime}$ and $\boldsymbol{\alpha}_{1}=(\alpha_{10},\boldsymbol{\alpha}_{11}^{\prime},\boldsymbol{\alpha}_{12}^{\prime})^{\prime}$ where $\boldsymbol{\alpha}_{11}=(\alpha_{111},...,\alpha_{11k_{1}})^{\prime}$ and $\boldsymbol{\alpha}_{12}=(\alpha_{121},...,\alpha_{12k_{3}})^{\prime}$. For the data generating process we let $W=D-\bar{\mathbf{Z}}_{1}^{\prime}\boldsymbol{\alpha}_{1}$ and assume $W|A=1\sim N(0,\sigma_{W}^{2})$. We define the exposure model for $A$ as
$\log\left\{\frac{P(A=1|\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})}{1-P(A=1|\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})}\right\}=\bar{\mathbf{Z}}_{2}^{\prime}\boldsymbol{\alpha}_{2}$ (3)
with $\bar{\mathbf{Z}}_{2}=(1,\mathbf{Z}_{2}^{\prime})^{\prime}$ and $\boldsymbol{\alpha}_{2}=(\alpha_{20},\boldsymbol{\alpha}_{21}^{\prime},\boldsymbol{\alpha}_{22}^{\prime})^{\prime}$ where $\boldsymbol{\alpha}_{21}=(\alpha_{211},...,\alpha_{21k_{2}})^{\prime}$ and $\boldsymbol{\alpha}_{22}=(\alpha_{221},...,\alpha_{22k_{3}})^{\prime}$.
2.3 A New Marginal Structural Model (MSM)
MSMs relate potential outcomes to exposures after marginalising over baseline covariates (Robins, 2000). We describe the effects of exposure status and dose through the MSM
$Y(A,D)=\psi_{0}+\psi_{11}A+\psi_{12}A(D-c)+Q$ (4)
where $Y(A,D)$ is the potential outcome, $(A,D)\perp Q$, $Q$ has a mean of zero, and $\boldsymbol{\psi}=(\psi_{0},\psi_{11},\psi_{12})^{\prime}$. Based on Assumption 3, the potential outcomes $Y(A,D)$ are independent of observed exposures $(A,D)$ given covariates $\mathbf{X}$ (Rosenbaum and Rubin, 1983) which implies $Q$ is conditionally independent of $A$ and $D$. Under Assumptions 1-4, $\psi_{11}$ and $\psi_{12}$ represent the causal effects of the exposure at the mean dose and the causal dose-response effect; see Appendix B for more derivation details.
2.4 Regression Adjustment with Two Propensity Scores
For binary exposures the propensity score (PS) is the conditional probability of exposure given covariates (Rosenbaum and Rubin, 1983), while for continuous exposures the generalised PS is the conditional exposure density given covariates (Imai and Van Dyk, 2004). As noted by Imai and Van Dyk (2004), any function of the confounders that indexes the exposure distribution can serve as a PS. In our setting with a binary exposure $A$ and a continuous dose $D$ among the exposed, we therefore define two propensity scores $\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})=\{S_{1}(\mathbf{Z}_{1};\boldsymbol{\alpha}_{1}),S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})\}^{\prime}$ with $\boldsymbol{\alpha}=(\boldsymbol{\alpha}_{1}^{\prime},\boldsymbol{\alpha}_{2}^{\prime})^{\prime}$ where $S_{1}(\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})=E(D|A=1,\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})$ and $S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})=E(A|\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})$.
To adjust for confounding, we consider the regression model
$E\{Y|A,D,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha});\boldsymbol{\eta}\}=\eta_{0}+\eta_{11}A+\eta_{12}A(D-c)+\eta_{21}S_{1}(\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})+\eta_{22}S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2}),$ (5)
where $\boldsymbol{\eta}=(\eta_{0},\boldsymbol{\eta}_{1}^{\prime},\boldsymbol{\eta}_{2}^{\prime})^{\prime}$ with $\boldsymbol{\eta}_{1}=(\eta_{11},\eta_{12})^{\prime}$ and $\boldsymbol{\eta}_{2}=(\eta_{21},\eta_{22})^{\prime}$. Under Assumptions 14 and correct specification of the exposure models, conditioning on $\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})$ balances covariates, and $\eta_{11}$ and $\eta_{12}$ identify the causal effect of exposure at the reference dose and the causal dose–response effect, respectively. Detailed derivations are given in Appendix C.
3 Two-stage Causal Analyses
While attractive in its simplicity, the regression adjustment strategy in Section 2.4 hinges on the correct specification of both propensity score models. This reliance on joint specification motivates a two-stage formulation that separates estimation of the dose–response from the contrast between exposure and no exposure at a reference dose. In Stage I we estimate the causal dose–response effect among exposed individuals, and in Stage II we estimate the effect of exposure at a reference dose. We use PS regression adjustment in Stage I, and PS regression adjustment, IPW, or AIPW in Stage II, and derive the corresponding estimating equations and asymptotic properties.
Let $n$ denote the sample size and index individuals by $i=1,\ldots,n$. The observed data are ${\mathcal{D}}_{i}=(Y_{i},\mathbf{X}_{i},A_{i},D_{i})$, $i=1,\ldots,n$. We write $S_{iq}=S_{q}(\mathbf{Z}_{iq};\boldsymbol{\alpha}_{q})$, $q=1,2$, for the propensity scores based on covariates $\mathbf{Z}_{i1}$ for $D\mid A=1$ and $\mathbf{Z}_{i2}$ for $A$, with parameters $\boldsymbol{\alpha}_{q}$.
3.1 Stage I: The Causal Dose-response Effect
Among exposed individuals ($A=1$), we consider a PS regression adjustment to assess the causal dose–response effect by fitting the model
$E\left\{Y_{i}|A_{i}=1,D_{i},S_{1}(\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1});\boldsymbol{\gamma}_{1}\right\}=\gamma_{10}+\gamma_{11}D_{i}+\gamma_{12}S_{1}(\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1})=\mu_{i11}(\boldsymbol{\phi}_{1}),$ (6)
by least squares where $S_{1}(\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1})=E(D_{i}|A_{i}=1,\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1})$, $\boldsymbol{\gamma}_{1}=(\gamma_{10},\gamma_{11},\gamma_{12})^{\prime}$, and $\boldsymbol{\phi}_{1}=(\boldsymbol{\gamma}_{1}^{\prime},\boldsymbol{\alpha}_{1}^{\prime})^{\prime}$. The estimating equation for $\boldsymbol{\gamma}_{1}$ is
$\mathbf{U}_{11}(\mathcal{D};\boldsymbol{\phi}_{1})=\sum_{i=1}^{n}\mathbf{U}_{i11}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})=\mathbf{0},$ (7)
where
$\mathbf{U}_{i11}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})=A_{i}\frac{\partial\mu_{i11}(\boldsymbol{\phi}_{1})}{\partial\boldsymbol{\gamma}_{1}}\{Y_{i}-\mu_{i11}(\boldsymbol{\phi}_{1})\}.$
The parameter $\boldsymbol{\alpha}_{1}$ can be estimated by solving
$\mathbf{U}_{12}(\mathcal{D};\boldsymbol{\alpha}_{1})=\sum^{n}_{i=1}\mathbf{U}_{i12}(\mathcal{D}_{i};\boldsymbol{\alpha}_{1})=\mathbf{0}$ (8)
where
$\mathbf{U}_{i12}(\mathcal{D}_{i};\boldsymbol{\alpha}_{1})=A_{i}\frac{\partial\mu_{i12}(\boldsymbol{\alpha}_{1})}{\partial\boldsymbol{\alpha}_{1}}\left\{D_{i}-\mu_{i12}(\boldsymbol{\alpha}_{1})\right\}$
with $\mu_{i12}(\boldsymbol{\alpha}_{1})=\bar{\mathbf{Z}}_{i1}^{\prime}\boldsymbol{\alpha}_{1}$. Considering (7) and (8) jointly, the Stage I estimating equation is
$\mathbf{U}_{1}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})=\sum_{i=1}^{n}\mathbf{U}_{i1}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})=\mathbf{0}$ (9)
with $\mathbf{U}_{i1}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})=\left\{\mathbf{U}_{i11}^{\prime}(\mathcal{D}_{i};\boldsymbol{\phi}_{1}),\mathbf{U}_{i12}^{\prime}(\mathcal{D}_{i};\boldsymbol{\alpha}_{1})\right\}^{\prime}.$ Under Assumptions 1–4, the causal dose–response effect is represented by $\gamma_{11}$, which corresponds to $\psi_{12}$ in (4).
3.2 Stage II: The Effect of Exposure Status
In Stage II, we assess the causal effect of exposure at a pre-specified reference dose $c$ compared to no exposure, with ${\gamma}_{11}$ obtained from Stage I and incorporated through the offset term ${\gamma}_{11}A_{i}(D_{i}-c)$. We use the term “offset" because in practice $\gamma_{11}$ will be estimated in Stage I and $\hat{\gamma}_{11}A_{i}(D_{i}-c)$ will be treated as a fixed quantity in what follows. We represent it as the scaled parameter $\gamma_{11}A_{i}(D_{i}-c)$ here to unify the discussion of the estimation. We consider three alternative approaches to estimation: (i) PS regression adjustment (Vansteelandt and Daniel, 2014), (ii) IPW (Robins, 2000), and (iii) AIPW (Bang and Robins, 2005; Robins et al., 2007).
3.2.1 Propensity score (PS) regression adjustment
To assess the effect of $A=1$ at dose $D=c$ via a PS regression adjustment, we consider a Stage II linear predictor of the form
$E\left\{Y_{i}|D_{i},A_{i},S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2});\boldsymbol{\gamma}_{2}\right\}=\mu_{i21}(\boldsymbol{\phi}_{2})+\text{offset}\left\{{\gamma}_{11}A_{i}(D_{i}-c)\right\}$
where $\mu_{i21}(\boldsymbol{\phi}_{2})=\gamma_{20}+\gamma_{21}A_{i}+\gamma_{22}S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})$ with PS $S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})=E(A_{i}|\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})$, and $\boldsymbol{\gamma}_{2}=(\gamma_{20},\gamma_{21},\gamma_{22})^{\prime}$ with $\boldsymbol{\phi}_{2}=(\boldsymbol{\gamma}_{2}^{\prime},\boldsymbol{\alpha}_{2}^{\prime})^{\prime}$.
The regression parameter $\boldsymbol{\gamma}_{2}$ can be estimated for specified $(\boldsymbol{\gamma}_{1}^{\prime},\boldsymbol{\alpha}_{2}^{\prime})^{\prime}$ by the estimating equation
$\tilde{\mathbf{U}}_{21}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\sum^{n}_{i=1}\tilde{\mathbf{U}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\boldsymbol{0}$ (10)
where
$\tilde{\mathbf{U}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\frac{\partial{\mu}_{i21}(\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\gamma}_{2}}\big[Y_{i}-\left[{\mu}_{i21}(\boldsymbol{\phi}_{2})+\text{offset}\left\{{\gamma}_{11}A_{i}(D_{i}-c)\right\}\right]\big].$
The parameter $\boldsymbol{\alpha}_{2}$ can be estimated by solving
$\mathbf{U}_{22}\left(\mathcal{D};\boldsymbol{\alpha}_{2}\right)=\sum^{n}_{i=1}\mathbf{U}_{i22}\left(\mathcal{D}_{i};\boldsymbol{\alpha}_{2}\right)=\mathbf{0}$ (11)
where
$\mathbf{U}_{i22}\left(\mathcal{D}_{i};\boldsymbol{\alpha}_{2}\right)=\frac{\partial S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})}{\partial\boldsymbol{\alpha}_{2}}\frac{1}{S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})\left\{1-S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})\right\}}\left\{A_{i}-S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})\right\}.$
Any binary regression model can be used; we employ a logistic link. Let
$\tilde{\mathbf{U}}_{2}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\sum_{i=1}^{n}\tilde{\mathbf{U}}_{i2}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\mathbf{0}$
denote the joint Stage II estimating equation under PS regression adjustment with $\tilde{\mathbf{U}}_{i2}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\big\{\tilde{\mathbf{U}}_{i21}^{\prime}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2}),\mathbf{U}_{i22}^{\prime}(\mathcal{D}_{i};\boldsymbol{\alpha}_{2})\big\}^{\prime}.$ Then taken together with the estimating equation of Stage I, the joint estimating equation for the two-stage analysis under PS regression adjustment is
$\tilde{\mathbf{U}}(\mathcal{D};\boldsymbol{\Omega})=\left\{\begin{array}[]{l}\mathbf{U}_{1}(\mathcal{D};\boldsymbol{\phi}_{1})\\ \tilde{\mathbf{U}}_{2}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})\end{array}\right\}=\boldsymbol{0}$ (12)
for $\boldsymbol{\Omega}=(\boldsymbol{\phi}_{1}^{\prime},\boldsymbol{\phi}_{2}^{\prime})^{\prime}$. Under Assumptions 1 to 4, $\gamma_{21}$ represents the causal effect of exposure at the reference dose compared to no exposure. and this holds likewise for the IPW and AIPW approaches described in Sections 3.2.2 and 3.2.3.
3.2.2 Inverse Probability Weighting (IPW)
We construct inverse probability weights based on $S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})$ to fit the Stage II marginal structural model
$E\left\{Y_{i}(A_{i},D_{i});\boldsymbol{\gamma}_{1},\boldsymbol{\gamma}_{2}\right\}=\gamma_{20}+\gamma_{21}A_{i}+\text{offset}\left\{{\gamma}_{11}A_{i}(D_{i}-c)\right\}.$
An IPW estimating equation (Hernán et al., 2000; Robins, 2000) is
$\bar{\mathbf{U}}_{21}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\sum^{n}_{i=1}\bar{\mathbf{U}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\mathbf{0}$ (13)
where
$\bar{\mathbf{U}}_{i21}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\sum_{a=0}^{1}w_{i}(a;\boldsymbol{\alpha}_{2})\frac{\partial{\mu}_{21}(a;\boldsymbol{\gamma}_{2})}{\partial\boldsymbol{\gamma}_{2}}\big[Y_{i}-\left[{\mu}_{21}(a;\boldsymbol{\gamma}_{2})+\text{offset}\left\{{\gamma}_{11}A_{i}(D_{i}-c)\right\}\right]\big]$
where
${\mu}_{21}(a;\boldsymbol{\gamma}_{2})=\gamma_{20}+\gamma_{21}a,$ (14)
and
$w_{i}(a;\boldsymbol{\alpha}_{2})=\frac{I(A_{i}=a)}{S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})^{a}\left\{1-S_{2}(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})\right\}^{1-a}}$ (15)
for $a=0,1$. The joint Stage II estimating equation under IPW is
$\bar{\mathbf{U}}_{2}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\sum_{i=1}^{n}\bar{\mathbf{U}}_{i2}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\mathbf{0}$
where $\bar{\mathbf{U}}_{i2}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\left\{\bar{\mathbf{U}}_{i21}^{\prime}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2}),\mathbf{U}_{i22}^{\prime}(\mathcal{D}_{i};\boldsymbol{\alpha}_{2})\right\}^{\prime}$. The two-stage estimating equations with IPW are therefore given by
$\bar{\mathbf{U}}(\boldsymbol{\Omega})=\left\{\begin{array}[]{l}\mathbf{U}_{1}(\mathcal{D};\boldsymbol{\phi}_{1})\\ \bar{\mathbf{U}}_{2}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})\end{array}\right\}=\mathbf{0}.$ (16)
3.2.3 Augmented Inverse Probability Weighting (AIPW)
An AIPW estimating function is introduced by Bang and Robins (2005) to obtain consistent estimators of causal effects when at least one of the PS model and the imputation model is correctly specified (Funk et al., 2011); the augmentation term we describe ensures this double robustness property. We again estimate ${\boldsymbol{\gamma}}_{2}$ via an augmented estimating equation
$\bar{\bar{\mathbf{U}}}_{21}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\sum^{n}_{i=1}\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\boldsymbol{0},$ (17)
and $\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})$ is given by
$\sum_{a=0}^{1}w_{i}(a;\boldsymbol{\alpha}_{2})\frac{\partial\mu_{12}(a;\boldsymbol{\gamma}_{2})}{\partial\boldsymbol{\gamma}_{2}}\big[Y_{i}-\left[\mu_{12}(a;\boldsymbol{\gamma}_{2})+\text{offset}\left\{{\gamma}_{11}A_{i}(D_{i}-c)\right\}\right]\big]-\{w_{i}(a;\boldsymbol{\alpha}_{2})-1\}\boldsymbol{g}_{i}(a;\boldsymbol{\theta},\boldsymbol{\gamma}_{2})$
where $\boldsymbol{\phi}_{2}=(\boldsymbol{\gamma}_{2}^{\prime},\boldsymbol{\alpha}_{2}^{\prime},\boldsymbol{\theta}^{\prime})^{\prime}$ with $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1}^{\prime},\boldsymbol{\theta}_{0}^{\prime})^{\prime}$, and $\mu_{12}(a;\boldsymbol{\gamma}_{2})$ is defined in (14). Here $w_{i}(a;\boldsymbol{\alpha}_{2})$ is given in (15), and
$\boldsymbol{g}_{i}(a;\boldsymbol{\theta},\boldsymbol{\gamma}_{2})=\frac{\partial\mu_{12}(a;\boldsymbol{\gamma}_{2})}{\partial\boldsymbol{\gamma}_{2}}\left\{m_{ia}(\boldsymbol{\theta})-\mu_{12}(a;\boldsymbol{\gamma}_{2})\right\}$
where $m_{ia}(\boldsymbol{\theta})$ denotes the imputed model based expression for the expected response under treatment $a\in{0,1}$; from (1). The imputed values are specified as
$m_{ia}(\boldsymbol{\theta}_{a})=E\left\{Y_{i}-\gamma_{11}a(D_{i}-c)\,|A_{i}=a,\mathbf{X}_{i};\boldsymbol{\theta}_{a}\right\}=\theta_{a0}+\boldsymbol{\theta}_{a1}^{\prime}\mathbf{X}_{i1}+\boldsymbol{\theta}_{a2}^{\prime}\mathbf{X}_{i2}+\boldsymbol{\theta}_{a3}^{\prime}\mathbf{X}_{i3},$
where $\boldsymbol{\theta}_{ap}$ is $k_{p}$-dimensional, $p=1,2,3$ for $a=0,1$. The estimating equation for $\boldsymbol{\theta}$ with specified $\boldsymbol{\gamma}_{1}$ is
$\mathbf{U}_{23}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\theta})=\sum^{n}_{i=1}\mathbf{U}_{i23}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta})=\mathbf{0}$ (18)
where $\mathbf{U}_{i23}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta})=\left\{\mathbf{U}_{i231}^{\prime}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta}_{1}),\mathbf{U}_{i232}^{\prime}(\mathcal{D}_{i};\boldsymbol{\theta}_{0})\right\}^{\prime}$ with
$\mathbf{U}_{i231}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta}_{1})=A_{i}\frac{\partial m_{i1}(\boldsymbol{\theta}_{1})}{\partial\boldsymbol{\theta}_{1}}\left\{Y_{i}-{\gamma}_{11}(D_{i}-c)-m_{i1}\left(\boldsymbol{\theta}_{1}\right)\right\},$
and
$\mathbf{U}_{i232}(\mathcal{D}_{i};\boldsymbol{\theta}_{0})=(1-A_{i})\frac{\partial m_{i0}(\boldsymbol{\theta}_{0})}{\partial\boldsymbol{\theta}_{0}}\left\{Y_{i}-m_{i0}(\boldsymbol{\theta}_{0})\right\}.$
By combining (17), (11), and (18), let $\bar{\bar{\mathbf{U}}}_{i2}(\mathcal{D}_{i};\boldsymbol{\phi}_{2})=\big\{\bar{\bar{\mathbf{U}}}_{i21}^{\prime}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2}),{{\mathbf{U}}}_{i22}^{\prime}(\mathcal{D}_{i};\boldsymbol{\alpha}_{2}),{{\mathbf{U}}}_{i23}^{\prime}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta})\big\}^{\prime}$ and consider the joint Stage II AIPW estimating equation
$\bar{\bar{\mathbf{U}}}_{2}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\sum_{i=1}^{n}\bar{\bar{\mathbf{U}}}_{i2}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})=\mathbf{0}.$ (19)
By combining (9) and (19), the joint estimating equation is
$\bar{\bar{\mathbf{U}}}(\boldsymbol{\Omega})=\left\{\begin{array}[]{l}\mathbf{U}_{1}(\mathcal{D};\boldsymbol{\phi}_{1})\\ \bar{\bar{\mathbf{U}}}_{2}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})\end{array}\right\}=\mathbf{0}.$ (20)
3.3 Estimation and Statistical Inference
3.3.1 Estimation in Two-stage Analysis
The two-stage analysis can be characterised by
$\mathbf{U}(\mathcal{D};\boldsymbol{\Omega})=\left\{\begin{array}[]{l}\mathbf{U}_{1}(\mathcal{D};\boldsymbol{\phi}_{1})\\ \mathbf{U}_{2}(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})\end{array}\right\}=\boldsymbol{0},$ (21)
where $\mathbf{U}_{1}(\cdot)$ is defined in (9) and $\mathbf{U}_{2}(\cdot)$ denotes the Stage II estimating equations under regression adjustment, IPW, or AIPW, as given in (12), (16), and (20), respectively. From a theoretical perspective, treating (21) as a unified system is convenient for deriving the large sample properties of the resulting estimators, and the estimates $\hat{\boldsymbol{\Omega}}=(\hat{\boldsymbol{\phi}}_{1}^{\prime},\hat{\boldsymbol{\phi}}_{2}^{\prime})^{\prime}$ can be obtained by solving (21) directly. In practice, however, estimation proceeds sequentially: the Stage I estimating equation (9) is solved to obtain $\hat{\boldsymbol{\phi}}_{1}$ and hence $\hat{\gamma}_{11}$, which is then substituted into the Stage II estimating equation. Solving the latter yields $\hat{\boldsymbol{\phi}}_{2}$ which includes $\hat{\gamma}_{21}$, the estimator of the causal effect of exposure at the reference dose.
3.3.2 Large Sample Theory
We now establish the large sample properties of the two-stage estimator with AIPW in Stage II; results for regression adjustment and IPW follow analogously. Let $\hat{\boldsymbol{\Omega}}$ denote the solution to the joint estimating equation (21), where $\mathbf{U}_{i}(\cdot)$ is the stacked vector of estimating functions contributions from individual $i$ from Stages I and II, $i=1,\ldots,n$.
Theorem 1
Suppose Assumptions 14 hold, and the PS $S_{1}(\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})=E(D|A=1,\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})$ is correctly specified. Then the solution $\hat{\boldsymbol{\phi}}_{1}=(\hat{\boldsymbol{\gamma}}_{1}^{\prime},\hat{\boldsymbol{\alpha}}_{1}^{\prime})^{\prime}$ to (9) is consistent for $\boldsymbol{\phi}_{1}$. In particular, $\hat{\gamma}_{11}$ is a consistent estimator of $\psi_{12}$, the causal dose–response effect in the marginal structural model (4).
Theorem 2
Suppose Assumptions 14 hold, and that at least one of the PS $S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})$ or the imputation model $m_{a}(\boldsymbol{\theta})$ is correctly specified. If $\hat{\gamma}_{11}$ is a consistent estimator of $\psi_{12}$, then the solution $\hat{\boldsymbol{\gamma}}_{2}=(\hat{\gamma}_{20},\hat{\gamma}_{21},\hat{\gamma}_{22})^{\prime}$ to the estimating equation (17) is consistent for $\boldsymbol{\gamma}_{2}$. In particular, $\hat{\gamma}_{21}$ is a consistent estimator of $\psi_{11}$, the causal effect of exposure at the reference dose in the marginal structural model (4).
Theorem 3
Suppose Assumptions 14 hold and standard regularity conditions (Van der Vaart, 2000) are satisfied. Then the joint estimator $\hat{\boldsymbol{\Omega}}$ is consistent for $\boldsymbol{\Omega}$, and $\sqrt{n}\,(\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega})$ converges in distribution to a mean-zero multivariate normal vector with covariance matrix
$\boldsymbol{\Sigma}(\boldsymbol{\Omega})=\mathcal{A}^{-1}(\boldsymbol{\Omega})\,\mathcal{B}(\boldsymbol{\Omega})\,\{\mathcal{A}^{-1}(\boldsymbol{\Omega})\}^{\prime},$
where $\mathcal{A}(\boldsymbol{\Omega})=E\left(-{\partial\mathbf{U}(\boldsymbol{\Omega})}/{\partial\boldsymbol{\Omega}^{\prime}}\right)$ and $\mathcal{B}(\boldsymbol{\Omega})=E\left\{\mathbf{U}_{i}(\boldsymbol{\Omega})\mathbf{U}_{i}^{\prime}(\boldsymbol{\Omega})\right\}$.
Proofs of Theorems 13 and explicit expressions for the sandwich covariance estimator and Wald statistics are given in Appendix D. We use the resulting sandwich covariance matrix for $\hat{\boldsymbol{\Omega}}$ to compute standard errors and construct Wald tests and confidence intervals.
4 Implication of Misspecified Propensity Scores in Two-stage Approaches
The consistency of the causal effect estimators introduced above relies on correct specification of the PS models. Akkaya Hocagil et al. (2021) derive the explicit form of the possibly biased estimator obtained by PS regression adjustment based on a possibly misspecified PS model for a semi-continuous exposure in one stage. Here we derive limiting values under misspecified propensity scores for the two-stage approach with PS regression adjustment in both stages.
4.1 Misspecification of the Stage I Propensity Score
We work with expectations of the estimating functions to characterise limiting values under misspecification (White, 1982). Since the data are i.i.d., we suppress the subscript $i$ and use $(Y,A,D,\mathbf{X})$ to denote a generic observation.
Let $\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})$ denote the Stage I PS under misspecification parameterized by $\tilde{\boldsymbol{\alpha}}_{1}$. A regression adjustment based on $\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})$ involves fitting
$E(Y|A=1,D,\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1});\tilde{\boldsymbol{\gamma}}_{1})=\tilde{\gamma}_{10}+\tilde{\gamma}_{11}D+\tilde{\gamma}_{12}\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})=\mu_{11}(\tilde{\boldsymbol{\phi}}_{1}),$
where $\tilde{\boldsymbol{\gamma}}_{1}=(\tilde{\gamma}_{10},\tilde{\gamma}_{11},\tilde{\gamma}_{12})^{\prime}$ and $\tilde{\boldsymbol{\phi}}_{1}=(\tilde{\boldsymbol{\gamma}}_{1}^{\prime},\tilde{\boldsymbol{\alpha}}_{1}^{\prime})^{\prime}$. Let $\mathcal{U}(\cdot)$ denote the corresponding estimating functions under misspecification. The limiting value of the estimator for $\boldsymbol{\gamma}_{1}$ is the solution to
$E\left\{\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})\right\}=\mathbf{0},$ (22)
with
$\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})=A\,\frac{\partial\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})}{\partial\tilde{\boldsymbol{\gamma}}_{1}}\{Y-\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})\},$
where the expectation is taken with respect to the true distribution of $(Y,A,D,\mathbf{X})$. Solving (22) yields $\boldsymbol{\gamma}_{1}^{\ast}=(\gamma_{10}^{\ast},\gamma_{11}^{\ast},\gamma_{12}^{\ast})^{\prime}$ with
$\gamma_{11}^{\ast}=\theta_{12}+\frac{\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)-\boldsymbol{\beta}_{1}(\mathbf{X}|A=1)\rho_{1}\sqrt{\tfrac{\rm var\{E(D|\mathbf{X},A=1)\}}{\rm var\{\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})\}}}\right]}{\rm var(D|A=1)-\rm var\{E(D|\mathbf{X},A=1)\}\rho_{1}^{2}},$ (23)
where $\rho_{1}=\text{corr}\{E(D|\mathbf{X},A=1),\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})\}$, $\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)=\{\zeta_{11}(\mathbf{X}|A=1),\ldots,\zeta_{1k}(\mathbf{X}|A=1)\}^{\prime}$, and $\boldsymbol{\beta}_{1}(\mathbf{X}|A=1)=\{\beta_{11}(\mathbf{X}|A=1),\ldots,\beta_{1k}(\mathbf{X}|A=1)\}^{\prime}$, with $\zeta_{1j}(\mathbf{X}|A=1)=\text{cov}\{E(D|\mathbf{X},A=1),\mathbf{X}_{j}|A=1\}$ and $\beta_{1j}(\mathbf{X}|A=1)=\text{cov}\{\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1}),\mathbf{X}_{j}|A=1\}$ for $j=1,\ldots,k$. Further details are provided in Appendix E.1.
When the PS model for the continuous exposure is correctly specified, $\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})=E(D|\mathbf{X},A=1)$. In this case $\rho_{1}=1$, $\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)=\boldsymbol{\beta}_{1}(\mathbf{X}|A=1)$, and $\rm var\{E(D|\mathbf{X},A=1)\}=\rm var\{\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})\}$. Substituting these equalities into (23) eliminates the second term, yielding $\gamma_{11}^{\ast}=\theta_{12}$. Otherwise, when $\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})$ is misspecified, the asymptotic bias of $\hat{\gamma}_{11}$ depends on the covariance structure among $E(D|\mathbf{X},A=1)$, $\mathbf{X}$ given $A=1$, and $\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})$.
4.2 Biased Estimator in Stage II
We now derive the limiting value of the exposure effect at the reference dose compared to no exposure. Let $\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$ be the Stage II PS under misspecification, and then fit the regression model by treating $\tilde{\gamma}_{11}A(D-c)$ as an offset term in
$E\left\{Y|D,A,\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2});\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\gamma}}_{2}\right\}=\tilde{\gamma}_{20}+\tilde{\gamma}_{21}A+\tilde{\gamma}_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})+\text{offset}\left\{\tilde{\gamma}_{11}A(D-c)\right\}$
where $\tilde{\boldsymbol{\gamma}}_{2}=(\tilde{\gamma}_{21},\tilde{\gamma}_{22},\tilde{\gamma}_{23})^{\prime}$. The estimating equation for $\tilde{\boldsymbol{\gamma}}_{2}$ is
$E\left\{\mathcal{U}_{21}(D;\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\phi}}_{2})\right\}=\mathbf{0}$ (24)
where
$\mathcal{U}_{21}(D;\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\phi}}_{2})=\frac{\partial\mu_{21}(\tilde{\boldsymbol{\phi}}_{2})}{\partial\tilde{\boldsymbol{\phi}}_{2}}\left[Y-\left\{\mu_{21}(\tilde{\boldsymbol{\phi}}_{2})+\tilde{\gamma}_{11}A(D-c)\right\}\right]$
with $\mu_{21}(\tilde{\boldsymbol{\phi}}_{2})=\tilde{\gamma}_{20}+\tilde{\gamma}_{21}A+\tilde{\gamma}_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$ again; the expectation in (24) is taken with respect to the true distribution of $(Y,D,A,\mathbf{X})$. Solving (24) with ${\gamma}_{11}^{*}$ given by (23) yields $\boldsymbol{\gamma}_{2}^{*}=({\gamma}_{20}^{*},{\gamma}_{21}^{*},{\gamma}_{22}^{*})^{\prime}$. Here we let $\rho_{21}=\text{corr}\bigl\{E(A|\mathbf{X}),\tilde{S}_{1}(\mathbf{X})\bigr\}$, $\rho_{22}=\text{corr}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{1}(\mathbf{X})\bigr\}$, $\boldsymbol{\zeta}_{2}(\mathbf{X})=\bigl\{\zeta_{21}(\mathbf{X}),...,\zeta_{2k}(\mathbf{X})\bigr\}^{\prime}$, $\boldsymbol{\beta}_{2}(\mathbf{X})=\left\{\beta_{21}(\mathbf{X}),...,\beta_{2k}(\mathbf{X})\right\}^{\prime}$, and $\delta_{1}=\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{1}(\mathbf{X})\bigr\}$ where $\zeta_{2j}(\mathbf{X})=\text{cov}\left\{E(A|\mathbf{X}),\mathbf{X}_{j}\right\}$ and $\beta_{2j}(\mathbf{X})=\text{cov}\bigl\{\tilde{S}_{1}(\mathbf{X}),\mathbf{X}_{j}\bigr\}$ for j=1,2,,kj=1,2,...,k. Then ${\gamma}_{21}^{*}$ corresponding to the limiting value of the causal effect of the exposure at the reference dose $c$ compared to no exposure under misspecification is
$\gamma_{21}^{*}=\theta_{11}+\frac{\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{2}(\mathbf{X})-\boldsymbol{\beta}_{2}(\mathbf{X})\rho_{21}\sqrt{\frac{\rm var\left\{E(A|\mathbf{X})\right\}}{\rm var\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\}}}\right]}{\rm var(A)-\rm var\left\{E(A|\mathbf{X})\right\}\rho_{21}^{2}}$
$+\frac{(\theta_{12}-{\gamma}_{11}^{*})\left[\text{cov}(AD,A)-cE\left\{\rm var(A|\mathbf{X})\right\}-\delta_{1}\rho_{22}\sqrt{\rm var\left\{E(A|\mathbf{X})\right\}\rm var\left\{E(AD|\mathbf{X})\right\}}\right]}{\rm var(A)-\rm var\left\{E(A|\mathbf{X})\right\}\rho_{21}^{2}}.$ (25)
Further details are provided in Appendix E.2.
Note that $\gamma_{21}^{*}=\theta_{11}$ when the second and third terms in (4.2) are zero. If $\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$ is correctly specified so that $\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})=E(A\mid\mathbf{X})$, and the Stage I estimator is consistent (implying ${\gamma}_{11}^{*}=\theta_{12}$), these terms vanish and $\gamma_{21}^{*}=\theta_{11}$. More generally, the asymptotic bias in the Stage II estimator depends on the covariance and variance structure of $E(A\mid\mathbf{X})$, $\mathbf{X}$ and $\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$, as well as any misspecification of $\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})$ in Stage I. The bias decreases as $\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$ approaches $E(A\mid\mathbf{X})$ and as $\tilde{\gamma}_{11}$ approaches $\theta_{12}$.
5 Simulation Studies
5.1 Data Generation and Simulation Design
We consider two data generation processes.
5.1.1 Data Generation Model 1
Let $\mathbf{X}=(X_{1},\ldots,X_{6})^{\prime}$ follow a multivariate normal distribution with mean zero, unit variances and pairwise correlation $0.2$ with $\mathbf{X}_{1}=(X_{11},X_{12})^{\prime}$, $\mathbf{X}_{2}=(X_{21},X_{22})^{\prime}$, and $\mathbf{X}_{3}=(X_{31},X_{32})^{\prime}$. The binary exposure $A|\mathbf{Z}_{2}$ is generated from model (3) with $\bar{\mathbf{Z}}_{2}=(1,\mathbf{X}_{2}^{\prime},\mathbf{X}_{3}^{\prime})^{\prime}$, $\boldsymbol{\alpha}_{21}=\left(\log 1.2,\log 2\right)^{\prime}$ and $\boldsymbol{\alpha}_{22}=\left(\log 0.8,\log 1.3\right)^{\prime}$. We solve for $\alpha_{20}$ such that $P(A=1)=E_{\mathbf{Z}_{2}}\left\{P(A=1|\mathbf{Z}_{2})\right\}=0.25,0.5$ or $0.75$ which give values $\alpha_{20}=-1.249$, $0.0008$ and $1.249$. For those exposed, the continuous exposure $D$ is generated from model (2) with $\bar{\mathbf{Z}}_{1}=(1,\mathbf{X}_{1}^{\prime},\mathbf{X}_{3}^{\prime})^{\prime}$, $\alpha_{10}=0$, $\boldsymbol{\alpha}_{11}=(1,0.5)^{\prime}$, $\boldsymbol{\alpha}_{12}=(0.6,0.8)^{\prime}$, and $W\sim N(0,1)$. If $A_{i}=0$ for each individual $D_{i}$ undefined. Finally, we generate the response variable $Y$ from model (1) with $\theta_{0}=0$, ${\theta}_{11}=4$, $\theta_{12}=0.5$, $\boldsymbol{\theta}_{21}=(0.6,0.8)^{\prime}$, $\boldsymbol{\theta}_{22}=(0.8,0.4)^{\prime}$, $\boldsymbol{\theta}_{23}=(0.3,0.7)^{\prime}$, and $E_{i}\sim N(0,1)$.
5.1.2 Data Generation Model 2
Let $\mathbf{X}=(X_{1},X_{2})^{\prime}$ follow a bivariate normal distribution with mean zero, unit variances and correlation $0.2$. We generate $A$ based on
$\log\left\{\frac{\pi(\mathbf{X})}{1-\pi(\mathbf{X})}\right\}=\alpha_{10}+\boldsymbol{\alpha}_{11}^{\prime}\mathbf{X}$ (26)
with $\boldsymbol{\alpha}_{11}=(\log 1.2,\log 2)^{\prime}$; we find $\alpha_{10}$ to give marginal probabilities of $A=1$ as $0.25,0.5$ and $0.75$ yielding $\alpha_{10}=-1.224$, $0.005$ and $1.236$. If $A=1$, the continuous exposure $D$ is generated from
$D=\alpha_{20}+\boldsymbol{\alpha}_{21}^{\prime}\mathbf{X}+W,$ (27)
where $\alpha_{20}=0$, $\boldsymbol{\alpha}_{21}=(1,0.5)^{\prime}$, $W\perp\mathbf{X}$, and $W\sim N(0,1)$. The second data generation model is the special case of (1)
$Y=\theta_{0}+{\theta}_{11}A+\theta_{12}A(D-c)+\mathbf{X}^{\prime}\boldsymbol{\theta}_{2}+E,$ (28)
with $\theta_{0}=0$, ${\theta}_{11}=4$, $\theta_{12}=0.5$, $\boldsymbol{\theta}_{2}=(0.8,0.4)^{\prime}$, and $E\sim N(0,1)$.
5.2 Methods
For each data generation model, we generate $2000$ replicated datasets of size $n=1000$ and estimate the causal parameters using six methods: (i) a naive least-squares fit of the MSM (4); (ii) the correctly specified response model; (iii) two–PS regression adjustment in a single stage; (iv) two-stage approach with PS regression; (v) two-stage approach with IPW; and (vi) two-stage approach with AIPW.
Brookhart et al. (2006) note that including covariates unrelated to the exposure but related to the outcome can reduce the variance of causal effect estimators without introducing bias. Accordingly, we consider two propensity score specifications: (i) including only exposure-related covariates and (ii) including all outcome-related covariates. In data-generation model 2 these sets coincide, so the two specifications are identical.
To examine robustness, we consider five misspecification scenarios: (i) misspecified Stage I PS $S_{1}$; (ii) misspecified Stage II PS $S_{2}$; (iii) misspecified imputation model $m_{a}(\mathbf{X};\boldsymbol{\theta})$; (iv) misspecification of $S_{1}$ and $m_{a}(\mathbf{X};\boldsymbol{\theta})$; and (v) misspecification of $S_{2}$ and $m_{a}(\mathbf{X};\boldsymbol{\theta})$. In all cases we focus on the setting with minimal propensity scores $\mathbf{S}=\{S_{1}(\mathbf{Z}_{1}),S_{2}(\mathbf{Z}_{2})\}^{\prime}$. Full details of how each model is misspecified are given in Appendix F.1.
5.3 Finite Sample Performance
Table 1: Empirical properties of estimators of causal effects with $P(A=1)=0.5$ under six methods of analysis based on correctly specified models.
Minimal PS: $\mathbf{S}=\{S_{1}(\mathbf{Z}_{1}),S_{2}(\mathbf{Z}_{2})\}^{\prime}$} Expanded PS: $\mathbf{S}=\{S_{1}(\mathbf{X}),S_{2}(\mathbf{X})\}^{\prime}$}
Method Exposure Effect Ebias ESE RSE ECP(%) Ebias ESE RSE ECP(%)
Data Generation Model 1
Naive $D$ $\theta_{12}$ 0.760 0.036 0.035 0.0 - - - -
AA $\theta_{11}$ 0.952 0.128 0.126 0.0 - - - -
DG model $D$ $\theta_{12}$ -0.001 0.028 0.027 94.2 - - - -
AA $\theta_{11}$ <0.001<0.001 0.067 0.068 95.2 - - - -
Two PS Reg $D$ $\theta_{12}$ -0.002 0.034 0.034 94.3 -0.002 0.033 0.032 95.0
AA $\theta_{11}$ <0.001<0.001 0.073 0.074 95.4 <0.001<0.001 0.067 0.068 95.0
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ -0.002 0.060 0.059 94.4 -0.002 0.046 0.045 94.2
AA $\theta_{11}$ -0.002 0.096 0.098 95.2 <0.001<0.001 0.068 0.069 95.6
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ -0.002 0.060 0.059 94.4 -0.002 0.046 0.045 94.2
AA $\theta_{11}$ -0.002 0.107 0.107 94.5 <0.001<0.001 0.082 0.083 95.5
$\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$ $D$ $\theta_{12}$ -0.002 0.060 0.059 94.4 -0.002 0.046 0.045 94.2
AA $\theta_{11}$ 0.001 0.070 0.071 95.5 <0.001<0.001 0.069 0.070 95.4
Data Generation Model 2
Naive $D$ $\theta_{12}$ 0.462 0.034 0.034 0.0 - - - -
AA $\theta_{11}$ 0.486 0.080 0.080 0.0 - - - -
DG model $D$ $\theta_{12}$ 0.001 0.035 0.034 94.4 - - - -
AA $\theta_{11}$ <0.001<0.001 0.068 0.068 95.4 - - - -
Two PS Reg $D$ $\theta_{12}$ 0.001 0.035 0.037 95.4 - - - -
AA $\theta_{11}$ <0.001<0.001 0.068 0.068 95.3 - - - -
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.002 0.046 0.045 93.9 - - - -
AA $\theta_{11}$ <0.001<0.001 0.069 0.069 95.1 - - - -
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.002 0.046 0.045 93.9 - - - -
AA $\theta_{11}$ <0.001<0.001 0.072 0.072 94.9 - - - -
$\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.002 0.046 0.045 93.9 - - - -
AA $\theta_{11}$ <0.001<0.001 0.070 0.070 94.4 - - - -
  • DG model: Data generation model; Two PS Reg: Two-dimensional PS regression adjustment; $\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$: Using PS regression adjustment in Stage I and II; $\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$: Using PS regression adjustment in Stage I and IPW in Stage II; $\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$: Using PS regression adjustment in Stage I and AIPW in Stage II.
Table 2: Empirical results for estimators of causal effects with $P(A=1)=0.5$ for two-propensity-score regression adjustment and two-stage analysis under model misspecification.
Misspecified $S_{1}$ Misspecified $S_{2}$
Method Exposure Effect Ebias ESE RSE ECP(%) Ebias ESE RSE ECP(%)
Data Generation Model 1
Two PS Reg $D$ $\theta_{12}$ 0.135 0.039 0.039 6.9 -0.002 0.035 0.036 96.0
AA $\theta_{11}$ 0.034 0.084 0.086 93.6 0.245 0.074 0.075 9.7
$\text{PS}_{\text{I}}$$\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.348 0.062 0.062 0.1 -0.002 0.060 0.059 94.4
AA $\theta_{11}$ 0.089 0.093 0.094 84.1 0.331 0.095 0.099 7.5
$\text{PS}_{\text{I}}$$\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.348 0.062 0.062 0.1 -0.002 0.060 0.059 94.4
AA $\theta_{11}$ 0.089 0.103 0.102 84.3 0.330 0.097 0.100 7.8
Data Generation Model 2
Two PS Reg $D$ $\theta_{12}$ -0.053 0.045 0.044 75.9 0.002 0.046 0.045 93.9
AA $\theta_{11}$ -0.016 0.073 0.072 94.3 0.208 0.070 0.070 15.8
$\text{PS}_{\text{I}}$$\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.144 0.044 0.043 8.4 0.002 0.046 0.045 93.9
AA $\theta_{11}$ 0.044 0.068 0.068 90.5 0.241 0.069 0.069 6.9
$\text{PS}_{\text{I}}$$\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.144 0.044 0.043 8.4 0.002 0.046 0.045 93.9
AA $\theta_{11}$ 0.043 0.072 0.071 90.3 0.241 0.070 0.069 6.8
  • $S_{1}$ represents the PS model for the continuous exposure $D$ conditioned on $A=1$, given by $S_{1}=E(D|A=1,\mathbf{Z}_{1})$; $S_{2}$ represents the PS model for the binary exposure $A$, given by $S_{2}=E(A|\mathbf{Z}_{2})$.
Table 3: Empirical results for estimators of effects for both parts with $P(A=1)=0.5$ for applying PS regression adjustment in Stage I and AIPW in Stage II under model misspecification.
Misspecified Model Exposure Effect Ebias ESE RSE ECP(%)
Data Generation Model 1
$S_{1}$ $D$ $\theta_{12}$ 0.348 0.063 0.062 0.1
AA $\theta_{11}$ 0.091 0.077 0.077 79.0
$S_{2}$ $D$ $\theta_{12}$ -0.002 0.060 0.059 94.4
AA $\theta_{11}$ -0.015 0.071 0.072 95.0
$m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ -0.002 0.060 0.059 94.4
AA $\theta_{11}$ -0.003 0.071 0.074 95.7
$S_{1}$ + $m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.348 0.062 0.062 0.1
AA $\theta_{11}$ 0.093 0.078 0.079 79.1
$S_{2}$ + $m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ -0.002 0.060 0.059 94.4
AA $\theta_{11}$ 0.225 0.069 0.071 11.0
Data Generation Model 2
$S_{1}$ $D$ $\theta_{12}$ 0.144 0.044 0.043 8.4
AA $\theta_{11}$ 0.044 0.070 0.069 90.8
$S_{2}$ $D$ $\theta_{12}$ 0.002 0.046 0.045 93.9
AA $\theta_{11}$ <0.001<0.001 0.069 0.069 95.0
$m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.002 0.046 0.045 93.9
AA $\theta_{11}$ <0.001<0.001 0.070 0.072 95.2
$S_{1}$ $m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.144 0.044 0.043 8.4
AA $\theta_{11}$ 0.044 0.070 0.070 91.0
$S_{2}$ + $m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.002 0.046 0.045 93.9
AA $\theta_{11}$ 0.241 0.070 0.069 6.8
  • $S_{1}$ represents the PS model for the continuous exposure $D$ conditioned on $A=1$, given by $S_{1}=E(D|A=1,\mathbf{Z}_{1})$; $S_{2}$ represents the PS model for the binary exposure $A$, given by $S_{2}=E(A|\mathbf{Z}_{2})$; $m_{a}(\mathbf{X};\boldsymbol{\theta})$ represents the imputation models for $a=0,1$.
Table 1 presents the empirical bias (EBias), empirical standard error (ESE), average robust standard error (RSE), and empirical coverage probability (ECP) for estimators with $P(A=1)=0.5$ under correctly specified models. The naive method yields poor estimates for both parts across both data-generation models. In contrast, all other methods show small biases and ECPs close to the nominal 95% level. Including all covariates in the PS models reduces empirical variance without inflating bias, in line with the variance reduction principle of Brookhart et al. (2006). The ESEs closely match the mean RSE, supporting the validity of the sandwich variance estimator in Section 3.3.2.
Table 2 reports empirical results under misspecified $S_{1}$ and $S_{2}$ for the two-dimensional PS regression adjustment, PS regression in both stages, and PS regression in Stage I with IPW in Stage II with $\pi^{*}=0.5$. Misspecification of $S_{1}$ induces bias in the Stage I estimator and consequently poor Stage II performance. Similarly, misspecification of $S_{2}$ yields large empirical bias in the Stage II estimator.
Table 3 summarises the performance of the Stage II AIPW estimators under model misspecification with $P(A=1)=0.5$. When either the imputation model or the $S_{2}$ model is misspecified, the estimator remains consistent by double robustness. In contrast, misspecified $S_{1}$ induces bias in both stages, and simultaneously misspecifying the imputation and $S_{2}$ models leads to biased Stage II estimates and loss of nominal coverage.
Tables 1–3 and 4–6 in Appendix F.2 present results for $P(A=1)=0.25$ and $0.75$ respectively, and show patterns consistent with those for $P(A=1)=0.5$.
6 Application to the prenatal alcohol exposure study
6.1 Description of the data
The Detroit Longitudinal Cohort is a prospective study of 480 pregnant African-American women from inner-city Detroit and their children, followed from birth to age 19 to investigate the effects of PAE (Jacobson et al., 2004, 2002). Prenatal alcohol intake was assessed at each visit using a timeline follow-back interview (Jacobson et al., 2002) and summarised as average daily consumption over pregnancy; let $T$ denote the average ounces of absolute alcohol (AA) per day. We treat PAE as a semi-continuous exposure: $A=I(T>0)$ is a binary indicator of any drinking, and $D=\log T$ is the log daily dose among drinkers. The reference dose is set to 2.31-2.31, the sample mean of $D$ among exposed mothers. The outcome is the Freedom from Distractibility Index of the Wechsler Intelligence Scales for Children, 3rd edition (WISC-III), measured at age 7 years for 377 children.
We adjust for baseline maternal sociodemographic characteristics, smoking and other substance use during pregnancy, reproductive history, gestational age at screening, and measures of the home environment and maternal cognition (listed in Table 7 in Appendix G.1). To address missingness, we generate 20 multiply imputed datasets using an imputation model including all variables; further details are given in Appendix G.2. Point estimates and standard errors are combined using Rubin’s rules under a missing-at-random assumption (Rubin, 1976; Little and Rubin, 2019).
6.2 Application and findings
Table 4: Estimated effects of drinking status and log prenatal absolute alcohol consumption per day
$D|A=1$ AA
Method Estimate Standard error Estimate Standard error
Covariate adjustment -0.125 0.515 0.253 1.655
Two–PS regression -0.189 0.572 -0.128 1.621
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ -0.011 0.518 -0.013 1.634
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ -0.268 2.128
$\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$ -0.176 1.965
  • Two–PS regression: two-dimensional PS regression adjustment; $\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$: PS regression adjustment in Stages I and II; $\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$: PS regression in Stage I and IPW in Stage II; $\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$: PS regression in Stage I and AIPW in Stage II.
We estimate the causal effects of prenatal drinking status and continuous dose using five approaches: (i) conventional covariate regression adjustment; (ii) two–propensity-score regression adjustment; (iii) a two-stage analysis with PS regression adjustment in Stage II; (iv) a two-stage analysis with IPW in Stage II; and (v) a two-stage analysis with AIPW in Stage II. Implementation details are provided in Appendix G.3.
Table 7 in Appendix G.1 summarises the fitted models for the continuous and binary exposure components, $E(D|A=1,\mathbf{X})$ and $P(A=1|\mathbf{X})$. Both models include all candidate covariates, following the recommendation of Brookhart et al. (2006) that including variables related to the outcome can improve precision of causal effect estimates.
Table 4 reports the estimated causal effects under the five approaches. For all PS–based methods, the estimated effects of both $\log(\text{AA/day})$ and drinking status are negative, with the magnitude of the drinking status effect larger than that of the dose effect. By contrast, the conventional covariate adjustment model yields a small positive estimate (0.253) for drinking status and a slightly negative estimate (0.125-0.125) for log dose. Given the reliance of covariate adjustment on correct specification of the outcome model (Greenland et al., 1999; Brookhart et al., 2006), these estimates may be more vulnerable to bias than those from the PS–based procedures.
For the log-dose effect among drinkers, PS–based point estimates are small and similar across methods, with standard errors around 0.5. For the drinking status effect, the PS–based estimates range from approximately 0.013-0.013 to 0.268-0.268, whereas the conventional covariate adjustment estimate is slightly positive. The two-stage AIPW estimator provides a negative point estimate (0.176-0.176) and benefits from its double-robustness property, but, as in our simulations, weighting-based methods yield larger standard errors than regression-based approaches (approximately 2.0 versus 1.6). Taken together, the PS–based results are compatible with modest harmful causal effects of both increased prenatal alcohol dose among drinkers and drinking status at the reference dose on WISC-III distractibility scores, although none of the estimated effects is statistically significant at the 5% level.
7 Discussion
We propose a causal framework involving a semi-continuous exposure with two-stage estimation. In Stage I, propensity score regression adjustment is used to estimate the continuous exposure effect, and Stage II focuses on the binary exposure via an AIPW estimator. We establish large sample properties for these estimators, characterise their limiting values under PS misspecification, and examine their finite sample performance. Together, these developments provide an interpretable framework that separates the consequences of any exposure from those of increasing dose, while allowing the use of doubly robust methods for the status effect. Theoretical results and simulations indicate that the proposed estimators perform well in finite samples and that the bias expressions under misspecification offer practical guidance for propensity score modelling with semi-continuous exposures.
A main limitation is that we do not pursue doubly robust estimation of the continuous exposure effect in Stage I. Doubly robust estimators for continuous treatments based on kernel smoothing and flexible generalised PS models (e.g. Zhu et al., 2015; Kennedy et al., 2017; Zhao et al., 2020) could be used to strengthen Stage I and reduce sensitivity to model misspecification.
Children’s cognition in the wider PAE literature is often assessed through multidimensional neurological outcomes across several U.S. cohort studies, including analyses based on hierarchical and structural equation models with PS adjustment (Akkaya Hocagil et al., 2022; Dang et al., 2023). Extending our semi-continuous two-stage framework to multivariate neurological outcomes, and to settings that pool information across cohorts, is a natural next step.
Acknowledgements
We thank Neil Dodge, Ph.D., for his assistance with data management support. This research was funded by grants to Sandra W. Jacobson and Joseph L. Jacobson from the National Institutes of Health/National Institute on Alcohol Abuse and Alcoholism (NIH/NIAAA; R01-AA025905) and the Lycaki-Young Fund from the State of Michigan. Richard J. Cook was supported by a grant from the Canadian Institutes for Health Research (PJT: 180551). Yeying Zhu is supported by a Discovery grant from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2017-04064). Louise Ryan was supported by the Australian Research Council Centre of Excellence for Mathematical and Statistical Frontiers (ACEMS) CE140100049. Data collection for the Detroit Longitudinal Study was supported by grants from NIH/NIAAA (R01-AA06966, R01-AA09524, and P50-AA07606) and NIH/ National Institute on Drug Abuse (R21- DA021034).
Appendix A. Identifying Assumptions
Assumption 5 (Stable Unit Treatment Value Assumption)
This assumption states that the potential outcome for one individual is independent of the exposure of another individual.
Assumption 6 (Consistency)
With a semi-continuous exposure, consistency is defined as $Y_{i}(1,d)=Y_{i}$ if the $i$-th individual is exposed at dose $D_{i}=d$ for any positive $d$, and $Y_{i}(0,d)=Y_{i}$ if unexposed. It assumes that the potential outcome of an individual under his/her observed exposure is exactly the outcome that is observed.
Assumption 7 (Ignorability)
With a semi-continuous exposure, the potential outcomes are independent of the observed exposure given covariates,
$Y_{i}(a,d)\perp A_{i},D_{i}|\mathbf{X}_{i}=\mathbf{x}.$
Specifically, the potential outcomes are independent of the observed continuous exposure given the $i$-th individual is exposed with dose $d$:
$Y_{i}(1,d)\perp D_{i}|A_{i}=1,\mathbf{X}_{i}=\mathbf{x}.$
Moreover, the potential outcome if the $i$-th individual is exposed at dose $d$ and the potential outcome if the individual is unexposed are independent of the observed exposure status given the covariates:
$Y_{i}(1,d),Y_{i}(0,0)\perp A_{i}|\mathbf{X}_{i}.$
Assumption 8 (Positivity)
The density $f(D_{i}=d|A_{i}=1,\mathbf{Z}_{i1})>0$ for all $d$ for those exposed, and $P(A_{i}=a|\mathbf{Z}_{i2})>0$ for $a=0,1$. This states that any individual in the exposed group ($A_{i}=1$) has a positive probability $f(d|A_{i}=1,\mathbf{Z}_{i1})\Delta d$ of a dose over $(d,d+\Delta)$ for all values of $d$, and every individual has a positive probability of either exposure level $A=0$ or $1$ given their covariates.
Appendix B. Derivation of the Marginal Structural Model Parameters
To connect the MSM (4) in the main text with the true data generation model (1) in the main text, note that
$E(Y|A,D)=E_{\mathbf{X}|A,D}\{E(Y|A,D,\mathbf{X})\}=\theta_{0}+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X}|A,D)+\theta_{11}A+\theta_{12}A(D-c).$
Thus, the parameters in the MSM (4) can be expressed in terms of those in the data generation model (1) as $\psi_{0}=\theta_{0}+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X}|A,D)$, $\psi_{11}=\theta_{11}$, and $\psi_{12}=\theta_{12}$.
Appendix C. Proof of Balancing for Regression Adjustment with Two Propensity Scores
We consider the data generation model
$Y=\theta_{0}+\theta_{11}A+\theta_{12}A(D-c)+\boldsymbol{\theta}_{21}^{\prime}\mathbf{X}_{1}+\boldsymbol{\theta}_{22}^{\prime}\mathbf{X}_{2}+\boldsymbol{\theta}_{23}^{\prime}\mathbf{X}_{3}+E,$ (C.1)
where $\boldsymbol{\theta}_{2}=(\boldsymbol{\theta}_{21}^{\prime},\boldsymbol{\theta}_{22}^{\prime},\boldsymbol{\theta}_{23}^{\prime})^{\prime}$, $\mathbf{X}=(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime},\mathbf{X}_{3}^{\prime})^{\prime}$, and $E$ has mean zero and is independent of $(A,D,\mathbf{X})$.
Define $\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})=({S}_{1}(\mathbf{Z}_{1};\boldsymbol{\alpha}_{1}),{S}_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2}))^{\prime}$ where $S_{1}(\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})=E(A|\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})$, and $S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})=E(D|A=1,\mathbf{Z}_{2};\boldsymbol{\alpha}_{2}).$ Our objective is to show that regression adjustment on the two–dimensional propensity score $\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})$ suffices to identify the target causal effects.
C.1 Dose–response Effect among the Exposed
Consider the effect of increasing the dose $D$ by one unit, conditional on exposure ($A=1$). Under the data generation model (C.1), this contrast is
$E\left\{Y|A=1,D=d+1,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}-E\left\{Y|A=1,D=d,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}$
$=$ $\left[\theta_{0}+\theta_{11}+\theta_{12}(d+1-c)+\boldsymbol{\theta}_{2}^{\prime}E\left\{\mathbf{X}|A=1,D=d+1,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}\right]$
$\quad-\left[\theta_{0}+\theta_{11}+\theta_{12}(d-c)+\boldsymbol{\theta}_{2}^{\prime}E\left\{\mathbf{X}|A=1,D=d,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}\right]$
$=$ $\theta_{12}+\boldsymbol{\theta}_{2}^{\prime}\left[E\left\{\mathbf{X}|A=1,D=d+1,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}-E\left\{\mathbf{X}|A=1,D=d,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}\right].$ (C.2)
To recover $\theta_{12}$, the term in square brackets in (C.2) must equal zero, i.e.,
$E\left\{\mathbf{X}|A=1,D=d+1,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}=E\left\{\mathbf{X}|A=1,D=d,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}.$ (C.3)
To see why this suffices, note that
$E\left\{\mathbf{X}|A=1,D,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}=E\left\{\mathbf{X}|A=1,D=S_{1}(\mathbf{Z}_{1};\boldsymbol{\alpha}_{1})+W,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\},$
where $W$ denotes the error term in the exposure model for $D|A=1,\mathbf{X}$ with $W\sim N(0,\sigma_{W}^{2})$. Since $W$ is independent of $\mathbf{X}$ given $(A=1,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha}))$, we obtain
$E\left\{\mathbf{X}|A=1,D,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}=E\left\{\mathbf{X}|A=1,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\},$
which is no longer a function of $D$.
Hence, condition (C.3) holds when $\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})$ is defined as above, and the contrast in (C.2) reduces to $\theta_{12}$. This corresponds to $\psi_{12}$, the causal dose–response effect in the marginal structural model (2.4) in the main text.
C.2 Effect of Exposure at the Reference Dose versus No Exposure
The contrast of interest is the effect of exposure at the reference dose $D=c$ compared with no exposure ($A=0$). Under the data-generating model (C.1),
$E\left\{Y|A=1,D=c,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}-E\left\{Y|A=0,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}$
$\quad=\left[\theta_{0}+\theta_{11}+\boldsymbol{\theta}_{2}^{\prime}E\left\{\mathbf{X}|A=1,D=c,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}\right]-\left[\theta_{0}+\boldsymbol{\theta}_{2}^{\prime}E\left\{\mathbf{X}|A=0,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}\right]$
$\quad=\theta_{11}+\boldsymbol{\theta}_{2}^{\prime}\left[E\left\{\mathbf{X}|A=1,D=c,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}-E\left\{\mathbf{X}|A=0,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}\right].$ (C.4)
To isolate $\theta_{11}$ it suffices that
$E\left\{\mathbf{X}|A=1,D=c,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}=E\left\{\mathbf{X}|A=0,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}.$ (C.5)
Expanding both sides of (C.5) gives
$E\left\{\mathbf{X}|A=1,D=c,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}$ $=E\left\{\mathbf{X}|D=c,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}=E\left\{\mathbf{X}|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\},$ (C.6)
$E\left\{\mathbf{X}|A=0,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}$ $=\frac{E\left\{\mathbf{X}\,I(A=0)|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}}{P\{A=0|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\}}.$ (C.7)
The first equality in (C.6) holds since $D$ is observed only when $A=1$, and the second holds since $D=S_{1}(\boldsymbol{Z}_{1};\boldsymbol{\alpha}_{1})+W$ with $W\perp\mathbf{X}|\mathbf{S}(\boldsymbol{Z};\boldsymbol{\alpha})$. The ratio in (C.7) uses Bayes’ rule for conditional expectations.
To simplify (C.7), rewrite the numerator and denominator as
$E\left\{\mathbf{X}\,I(A=0)|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}$ $=E\left[\mathbf{X}\,E\left\{I(A=0)|\mathbf{X},\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right]$
$=E\left\{\mathbf{X}\,P(A=0|\mathbf{X})|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}$
$=E\left[\mathbf{X}\,\{1-S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})\}|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right]$
$=\{1-S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})\}\,E\left\{\mathbf{X}|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}$ (C.8)
$P\{A=0|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\}$ $=E\left\{1-S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}=1-S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})$ (C.9)
The first line in (C.8) uses the law of iterated expectations. Next, $E\{I(A=0)|\mathbf{X},\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\}=P(A=0|\mathbf{X})$, and by the model $P(A=1|\mathbf{X})=S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})$ we have $P(A=0|\mathbf{X})=1-S_{2}(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2})$.
Substituting (C.8) and (C.9) into (C.7) gives
$E\left\{\mathbf{X}|A=0,\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}=E\left\{\mathbf{X}|\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})\right\}.$ (C.10)
Combining (C.6) and (C.10) yields (C.5), so (C.4) reduces to $\theta_{11}$, which coincides with $\psi_{11}$, the causal effect of exposure at the reference dose versus no exposure in the marginal structural model (2.4) in the main text.
Appendix D. Additional Materials on Statistical Inference
D.1 Proof of Theorem 1
Under Assumptions 1 to 4, and the propensity score $S_{1}(\boldsymbol{Z}_{1};\boldsymbol{\alpha}_{1})$ is correctly specified. We aim to show $E_{\mathcal{D}_{i}}\bigl\{\mathbf{U}_{i11}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})\bigr\}=\mathbf{0}.$ Firstly, we take conditional expectation respect to $Y_{i}|D_{i},A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3}$ to obtain $A_{i}{\partial\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)}/{\partial\boldsymbol{\gamma}_{1}}\left\{E\left(Y_{i}|D_{i},A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3}\right)-\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)\right\}.$ Then we take the expectation with respect to $\mathbf{X}_{i1}|D_{i},A_{i},\mathbf{X}_{i2},\mathbf{X}_{i3}$ to obtain
$\int_{\mathbf{X}_{1}}A_{i}\frac{\partial\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)}{\partial\boldsymbol{\gamma}_{1}}\left\{E\left(Y_{i}|D_{i},A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3}\right)-\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)\right\}\partial G\left(\mathbf{X}_{i1}|D_{i},A_{i},\mathbf{X}_{i2},\mathbf{X}_{i3}\right)$
where $\partial G\left(\mathbf{X}_{i1}|D_{i},A_{i},\mathbf{X}_{i2},\mathbf{X}_{i3}\right)$ is given by
$\frac{\partial F(D_{i}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i2}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i3})}{\partial F(D_{i},A_{i},\mathbf{X}_{i2},\mathbf{X}_{i3})}\partial F(A_{i}|\mathbf{X}_{i1},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i3}|\mathbf{X}_{i1})\partial G(\mathbf{X}_{i1}).$
Upon substitution
$\int_{\mathbf{X}_{1}}A_{i}\frac{\partial\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)}{\partial\boldsymbol{\gamma}_{1}}\left\{E\left(Y_{i}|D_{i},A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3}\right)-\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)\right\}$
$\cdot\frac{\partial F(D_{i}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i2}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i3})\partial F(A_{i}|\mathbf{X}_{i1},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i3}|\mathbf{X}_{i1})\partial G(\mathbf{X}_{i1})}{\partial F(D_{i},A_{i},\mathbf{X}_{i2},\mathbf{X}_{i3})}$ (D.1)
$=\int_{\mathbf{X}_{1}}A_{i}\frac{\partial\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)}{\partial\boldsymbol{\gamma}_{1}}\frac{\partial F(D_{i}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i2}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i3})\partial F(A_{i}|\mathbf{X}_{i1},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i3}|\mathbf{X}_{i1})}{\partial F(D_{i},A_{i},\mathbf{X}_{i2},\mathbf{X}_{i3})}$
$\cdot\left\{E\left(Y_{i}|D_{i},A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3}\right)-\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)\right\}\partial G(\mathbf{X}_{i1})$ (D.2)
$=\int_{\mathbf{X}_{1}}A_{i}\frac{\partial\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)}{\partial\boldsymbol{\gamma}_{1}}\frac{\partial F(D_{i}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i2}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i3})\partial F(A_{i}|\mathbf{X}_{i1},\mathbf{X}_{i3})\partial G(\mathbf{X}_{i3}|\mathbf{X}_{i1})}{\partial F(D_{i},A_{i},\mathbf{X}_{i2},\mathbf{X}_{i3})}$
$\cdot\bigl[E\left(Y_{i}|D_{i},\mathbf{X}_{i2},\mathbf{X}_{i3}\right)-E\left\{Y_{i}|A_{i},D_{i},S_{1}\left(\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1}\right);\boldsymbol{\gamma}_{1}\right\}\bigr]$ (D.3)
$=\mathbf{0}.$
Note (D.2) is an algebraic re-expression of (D.1). The transition from (D.2) to (D.3) follows from the equality $\mu_{i11}\left(\boldsymbol{\phi}_{1}\right)=E\left\{Y_{i}|A_{i},D_{i},S_{1}\left(\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1}\right);\boldsymbol{\gamma}_{1}\right\}$, together with the law of iterated expectations, under which integrating out $\mathbf{X}_{i1}$ removes its conditional dependence from the first expectation term inside the curly brackets of (D.2). Then, under the assumption that the generalized propensity score model $S_{1}\left(\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1}\right)$ is correctly specified, conditioning on $S_{1}\left(\mathbf{Z}_{i1};\boldsymbol{\alpha}_{1}\right)$ controls for confounding by $\mathbf{X}_{i1}$ and $\mathbf{X}_{i3}$ in estimating the causal effect of $D_{i}$ on $Y_{i}$ among the exposed. As a result, two conditional expectations in the brackets in (D.3) are equal yielding the desired result. Therefore, the estimator $\hat{\gamma}_{11}$ is a consistent estimator of $\psi_{12}$ in the MSM (2.4) in the main text under Assumptions 1-4.
D.2 Proof of Theorem 2
Under Assumptions 1 to 4, and assuming $\hat{\gamma}_{11}$ obtained from Stage I is consistent for $\psi_{12}$, we show the AIPW estimating function is an unbiased estimating function to prove consistency when at least one of the propensity score model $S_{2}\left(\mathbf{Z}_{2};\boldsymbol{\alpha}_{2}\right)$ and the imputation model $m_{a}(\boldsymbol{\theta})$ are correctly specified.
Proving consistency is equivalent to showing
$E_{\mathcal{D}_{i}}\left\{\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})\right\}=\mathbf{0}.$
We first rewrite the estimating function $\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})$ as
$\sum_{a=0}^{1}w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)\frac{\partial\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)}{\partial\gamma_{2}}\left[Y_{i}-\left\{\mu_{12}\left(a;\gamma_{2}\right)+\gamma_{11}A_{i}\left(D_{i}-c\right)\right\}\right]-\left\{w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)-1\right\}\boldsymbol{g}_{i}\left(a;\boldsymbol{\theta},\boldsymbol{\gamma}_{2}\right)$ (D.4)
$=$ $\sum_{a=0}^{1}w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)\frac{\partial\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)}{\partial\gamma_{2}}\left\{\bar{Y}_{i}-\mu_{12}\left(a;\gamma_{2}\right)\right\}-\left\{w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)-1\right\}\boldsymbol{g}_{i}\left(a;\boldsymbol{\theta},\boldsymbol{\gamma}_{2}\right)$ (D.5)
where $\bar{Y}_{i}=\theta_{0}+\theta_{11}A_{i}+\boldsymbol{\theta}_{21}^{\prime}\mathbf{X}_{1}+\boldsymbol{\theta}_{22}^{\prime}\mathbf{X}_{2}+\boldsymbol{\theta}_{23}^{\prime}\mathbf{X}_{3}$. Under the assumption that $\hat{\gamma}_{11}$ is a consistent estimator of $\psi_{12}$ in the MSM, where $\psi_{12}=\theta_{12}$ in the data-generating model, the term $\hat{\gamma}_{11}A_{i}(D_{i}-c)$ captures the component of $Y_{i}$ that is causally attributable to the continuous exposure. Subtracting this term from $Y_{i}$ in (D.4) removes the causal dose-response effect, isolating the portion of the linear predictor that depends only on the binary exposure and the covariates. This yields (D.5). Therefore, the problem is reduced to showing that
$E_{\bar{Y}_{i},A_{i},\mathbf{X}_{i}}\left\{\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})\right\}=\mathbf{0},$
since (D.5) no longer depends on the continuous exposure $D_{i}$, and $\bar{Y}_{i}$ captures only the component of the outcome explained by $A_{i}$ and $\mathbf{X}_{i}$.
Here, we consider three scenarios seperately: 1) only the imputation model is correctly specified, 2) only the propensity score model is correctly specified, 3) both are correctly specified. First, we take the conditional expectation of $\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})$ with respect to $\bar{Y}_{i}|A_{i},\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3}$. Then we get
$\sum_{a=0}^{1}w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)\frac{\partial\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)}{\partial\boldsymbol{\gamma}_{2}}\left\{E(\bar{Y}_{i}|A_{i},\mathbf{X}_{i})-\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)\right\}$
$-\left\{w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)-1\right\}\frac{\partial\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)}{\partial\boldsymbol{\gamma}_{2}}\left\{m_{ia}(\boldsymbol{\theta})-\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)\right\}$ (D.6)
where $w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)={I\left(A_{i}=a\right)}/\left[{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)^{a}\left\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\right\}^{1-a}}\right].$ The first term in (D.2 Proof of Theorem 2) can be represented as
$\sum_{a=0}^{1}\frac{I\left(A_{i}=a\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)^{a}\left\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\right\}^{1-a}}\frac{\partial\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)}{\partial\boldsymbol{\gamma}_{2}}\left\{E(\bar{Y}_{i}|A_{i},\mathbf{X}_{i})-\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)\right\}$
$=$ $\sum_{a=0}^{1}\frac{I\left(A_{i}=a\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)^{a}\left\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\right\}^{1-a}}\left(\begin{matrix}1\\ a\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i},\mathbf{X}_{i})-\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)\right\}$
$=$ $\frac{I(A_{i}=0)}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i},\mathbf{X}_{i})-\mu_{12}\left(0;\boldsymbol{\gamma}_{2}\right)\right\}+\frac{I(A_{i}=1)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i},\mathbf{X}_{i})-\mu_{12}\left(1;\boldsymbol{\gamma}_{2}\right)\right\}$ (D.7)
$=$ $\frac{(1-A_{i})}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i},\mathbf{X}_{i})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}+\frac{A_{i}}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i},\mathbf{X}_{i})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$ (D.8)
Then we take conditional expectation to (D.8) with respect to $A_{i}|\mathbf{X}_{i}$ to obtain
$P(A_{i}=1|\mathbf{X}_{i})\biggl[\frac{(1-1)}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=1,\mathbf{X}_{i})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$
$+\frac{1}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=1,\mathbf{X}_{i})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}\biggl]+$
$P(A_{i}=0|\mathbf{X}_{i})\biggl[\frac{(1-0)}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=0,\mathbf{X}_{i})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$
$+\frac{0}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=0,\mathbf{X}_{i})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}\biggl]$ (D.9)
$=$ $P(A_{i}=1|\mathbf{X}_{i})\frac{1}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=1,\mathbf{X}_{i})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$
$+P(A_{i}=0|\mathbf{X}_{i})\frac{1}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=0,\mathbf{X}_{i})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$ (D.10)
$=$ $P(A_{i}=1|\mathbf{Z}_{i1})\frac{1}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left[E(\bar{Y}_{i}|A_{i}=1,\mathbf{X}_{i})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right]$
$+P(A_{i}=0|\mathbf{Z}_{i1})\frac{1}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=0,\mathbf{X}_{i})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$ (D.11)
$=$ $\frac{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=1,\mathbf{X}_{i})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$
$+\frac{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{E(\bar{Y}_{i}|A_{i}=0,\mathbf{X}_{i})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$ (D.12)
$=$ $\boldsymbol{0}$
Thus, we only need to show the second term in (D.2 Proof of Theorem 2) is equal to $\boldsymbol{0}$ under the following three cases.
Case 1: only the imputation model is correctly specified.
Let $\tilde{S}_{2}\left(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}\right)$ denote the Stage II propensity score undermisspecification, and the weight corresponding to the misspecified propensity score is denoted as $\tilde{w}_{i}\left(a;\boldsymbol{\alpha}_{2}\right)$.
The second component in (D.2 Proof of Theorem 2) is
$\sum_{a=0}^{1}\left(\tilde{w}_{i}\left(a;\boldsymbol{\alpha}_{2}\right)-1\right)\frac{\partial\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)}{\partial\boldsymbol{\gamma}_{2}}\left\{m_{ia}(\boldsymbol{\theta})-\mu_{12}\left(a;\boldsymbol{\gamma}_{2}\right)\right\}.$
Since the imputation model is correctly specified,
$m_{a}(\boldsymbol{\theta})=E(\bar{Y}_{i}|A=a,\mathbf{X}_{i1},\mathbf{X}_{i2},\mathbf{X}_{i3})=\mu_{12}(a;\boldsymbol{\gamma}_{2}).$ (D.13)
Under (D.13), (D.2 Proof of Theorem 2) equals 0. Therefore, we conclude that $\hat{\gamma}_{21}$ is consistent for $\psi_{11}$ in the MSM (2.4) if the propensity score model is misspecified and the imputation model is correctly specified, under Assumptions 1–4 and based on a consistent $\hat{\gamma}_{11}$.
Case 2: only the propensity score model is correctly specified.
[…]
Since $A_{i}\in\{0,1\}$, the indicator functions $I(A_{i}=1)$ and $I(A_{i}=0)$ can be equivalently expressed as $A_{i}$ and $1-A_{i}$ respectively. Therefore, in the subsequent derivation, (D.7) can be rewritten as (D.8), and (D.14) as (D.15). Since the propensity score model is correctly specified, $S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)=E(A_{i}|\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2})$.
The second component in (D.2 Proof of Theorem 2) can be re-expressed as
$\sum_{a=0}^{1}\left[\frac{I(A_{i}=a)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)^{a}\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\}^{1-a}}-1\right]\left(\begin{matrix}1\\ a\end{matrix}\right)\left\{\tilde{m}_{a}(\tilde{\boldsymbol{\theta}})-\mu_{12}(a;\boldsymbol{\gamma}_{2})\right\}$
$=$ $\frac{I(A_{i}=0)-\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\}}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{\tilde{m}_{0}(\tilde{\boldsymbol{\theta}})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$
$+\frac{I(A_{i}=1)-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{\tilde{m}_{1}(\tilde{\boldsymbol{\theta}})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$ (D.14)
$=$ $\frac{(1-A_{i})-\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\}}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{\tilde{m}_{0}(\tilde{\boldsymbol{\theta}})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$
$+\frac{A_{i}-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{\tilde{m}_{1}(\tilde{\boldsymbol{\theta}})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$ (D.15)
Now, we take the conditional expectation of (D.15) with respect to $A_{i}|\mathbf{X}_{i}$ to obtain
$\frac{\{1-E(A_{i}|\mathbf{X}_{i})\}-\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\}}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{\tilde{m}_{0}(\tilde{\boldsymbol{\theta}})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$
$+\frac{E(A_{i}|\mathbf{X}_{i})-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{\tilde{m}_{1}(\tilde{\boldsymbol{\theta}})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$ (D.16)
$=$ $\frac{\{1-E(A_{i}|\mathbf{Z}_{i1})\}-\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\}}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{\tilde{m}_{0}(\tilde{\boldsymbol{\theta}})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}$
$+\frac{E(A_{i}|\mathbf{Z}_{i1})-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{\tilde{m}_{1}(\tilde{\boldsymbol{\theta}})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$ (D.17)
$=$ $\frac{\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\}-\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\}}{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 0\end{matrix}\right)\left\{\tilde{m}_{0}(\tilde{\boldsymbol{\theta}})-\mu_{12}(0;\boldsymbol{\gamma}_{2})\right\}+$
$+\frac{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)}\left(\begin{matrix}1\\ 1\end{matrix}\right)\left\{\tilde{m}_{1}(\tilde{\boldsymbol{\theta}})-\mu_{12}(1;\boldsymbol{\gamma}_{2})\right\}$ (D.18)
$=$ $\boldsymbol{0}$
Therefore, (D.2 Proof of Theorem 2) equals 0 in this case. In conclusion, $\hat{\gamma}_{21}$ is consistent for $\psi_{11}$ in the MSM (2.4) if the imputation model is misspecified and the propensity score model is correctly specified, under Assumptions 1-4 and based on a consistent $\hat{\gamma}_{11}$.
Case 3: both the imputation model and the propensity score model are correctly specified.
This case represents a more strict scenario. Since Cases 1 and 2 have respectively established that the augmented estimating function remains unbiased when either model is correctly specified, it immediately follows that the estimating function remains unbiased when both are correct.
In this setting, the bias from both sources is simultaneously removed, and the estimating function consistently recovers the causal estimand. Therefore, Case 3 is a special instance where both conditions for double robustness are satisfied, and the consistency result follows directly from the arguments in Cases 1 and 2.
D.3 Proof of Theorem 3
Here we derive the asymptotic covariance matrix of the estimator obtained from two-stage analysis focusing on applying PS regression adjustment in stage I and use of AIPW in stage II. Recall that ${\hat{\boldsymbol{\Omega}}}=(\hat{\boldsymbol{\phi}}_{1}^{\prime},\hat{\boldsymbol{\phi}}_{2}^{\prime})^{\prime}$ is the solution of
$\bar{\bar{\mathbf{U}}}(\boldsymbol{\Omega})=\left\{\begin{array}[]{l}\mathbf{U}_{1}\left(\mathcal{D};\boldsymbol{\phi}_{1}\right)\\ \bar{\bar{\mathbf{U}}}_{2}\left(\mathcal{D};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2}\right)\end{array}\right\}=\mathbf{0}.$
where $\boldsymbol{\phi}_{1}=(\boldsymbol{\gamma}_{1}^{\prime},\boldsymbol{\alpha}_{1}^{\prime})^{\prime}$ and $\boldsymbol{\phi}_{2}=(\boldsymbol{\gamma}_{2}^{\prime},\boldsymbol{\alpha}_{2}^{\prime})^{\prime}$.
Since
$\bar{\bar{\mathbf{U}}}({\hat{\boldsymbol{\Omega}}})=\bar{\bar{\mathbf{U}}}(\boldsymbol{\Omega})+\frac{\partial\bar{\bar{\mathbf{U}}}(\boldsymbol{\Omega})}{\partial\boldsymbol{\Omega}^{\prime}}\left({\bar{\bar{\boldsymbol{\Omega}}}}-\boldsymbol{\Omega}\right)+o_{p}\left(\frac{1}{\sqrt{n}}\right)$
then we have
$\sqrt{n}(\bar{\bar{\boldsymbol{\Omega}}}-\boldsymbol{\Omega})=\left\{-\frac{1}{n}\frac{\partial\bar{\bar{\mathbf{U}}}(\boldsymbol{\Omega})}{\partial\boldsymbol{\Omega}^{\prime}}\right\}^{-1}\left\{\frac{1}{\sqrt{n}}\bar{\bar{\mathbf{U}}}(\boldsymbol{\Omega})\right\}+o_{p}(1)$
where $\frac{\partial\bar{\bar{\mathbf{U}}}(\boldsymbol{\Omega})}{\partial\boldsymbol{\Omega}^{\prime}}$ is
$\sum_{i=1}^{n}\left\{\begin{matrix}\frac{\partial\mathbf{U}_{i11}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})}{\partial\boldsymbol{\gamma}_{1}}&\frac{\partial\mathbf{U}_{i11}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})}{\partial\boldsymbol{\alpha}_{1}}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{0}&\frac{\partial\mathbf{U}_{i12}(\mathcal{D}_{i};\boldsymbol{\alpha}_{1})}{\partial\boldsymbol{\alpha}_{1}}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\gamma}_{1}}&\boldsymbol{0}&\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\gamma}_{2}}&\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\alpha}_{2}}&\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\theta}_{1}}&\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\theta}_{2}}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\frac{\partial\mathbf{U}_{i22}\left(\mathcal{D}_{i};\boldsymbol{\alpha}_{2}\right)}{\partial\boldsymbol{\alpha}_{2}}&\boldsymbol{0}&\boldsymbol{0}\\ \frac{\partial\mathbf{U}_{i231}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta}_{1})}{\partial\boldsymbol{\gamma}_{1}}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\frac{\partial\mathbf{U}_{i231}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta}_{1})}{\partial\boldsymbol{\theta}_{1}}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\frac{\partial\mathbf{U}_{i232}(\mathcal{D}_{i};\boldsymbol{\theta}_{0})}{\partial\boldsymbol{\theta}_{0}}\end{matrix}\right\}$
As $n\to\infty$, $\mathcal{A}$ converges to
$E\left\{-\frac{\partial{\bar{\bar{\mathbf{U}}}_{i}}(\boldsymbol{\Omega})}{\partial\boldsymbol{\Omega}^{\prime}}\right\}=\left\{\begin{matrix}\mathcal{A}_{11}&\mathcal{A}_{12}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{0}&\mathcal{A}_{22}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \mathcal{A}_{31}&\boldsymbol{0}&\mathcal{A}_{33}&\mathcal{A}_{34}&\mathcal{A}_{35}&\mathcal{A}_{36}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\mathcal{A}_{44}&\boldsymbol{0}&\boldsymbol{0}\\ \mathcal{A}_{51}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\mathcal{A}_{55}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\mathcal{A}_{66}\end{matrix}\right\}$
where
$\mathcal{A}_{11}=E\left\{-\frac{\partial\mathbf{U}_{i11}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})}{\partial\boldsymbol{\gamma}_{1}}\right\}\text{, }\mathcal{A}_{12}=E\left\{-\frac{\partial\mathbf{U}_{i11}(\mathcal{D}_{i};\boldsymbol{\phi}_{1})}{\partial\boldsymbol{\alpha}_{2}}\right\},$
$\mathcal{A}_{22}=E\left\{-\frac{\partial\mathbf{U}_{i12}(\boldsymbol{\boldsymbol{\alpha}_{2})}}{\partial\boldsymbol{\alpha}_{2}}\right\}\text{, }\mathcal{A}_{31}=E\left\{\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\gamma}_{1}}\right\},$
$\mathcal{A}_{34}=E\left\{\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\gamma}_{2}}\right\}\text{, }\mathcal{A}_{33}=E\left\{\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\alpha}_{1}}\right\},$
$\mathcal{A}_{35}=E\left\{\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\theta}_{1}}\right\}\text{, }\mathcal{A}_{36}=E\left\{\frac{\partial\bar{\bar{\mathbf{U}}}_{i21}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\phi}_{2})}{\partial\boldsymbol{\theta}_{2}}\right\},$
$\mathcal{A}_{44}=E\left\{\frac{\partial\mathbf{U}_{i22}\left(\mathcal{D}_{i};\boldsymbol{\alpha}_{2}\right)}{\partial\boldsymbol{\alpha}_{1}}\right\}\text{, }\mathcal{A}_{51}=E\left\{\frac{\partial\mathbf{U}_{i231}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta}_{1})}{\partial\boldsymbol{\gamma}_{1}}\right\},$
$\mathcal{A}_{55}=E\left\{\frac{\partial\mathbf{U}_{i231}(\mathcal{D}_{i};\boldsymbol{\gamma}_{1},\boldsymbol{\theta}_{1})}{\partial\boldsymbol{\theta}_{1}}\right\}\text{, and }\mathcal{A}_{66}=E\left\{\frac{\partial\mathbf{U}_{i232}(\mathcal{D}_{i};\boldsymbol{\theta}_{0})}{\partial\boldsymbol{\theta}_{0}}\right\}.$
As $n\to\infty$,
$\frac{1}{n}\sum_{i=1}^{n}{\bar{\bar{\mathbf{U}}}_{i}}(\boldsymbol{\Omega}){\bar{\bar{\mathbf{U}}}_{i}}(\boldsymbol{\Omega})^{\prime}$
converges in probability to
$E\left\{{\bar{\bar{\mathbf{U}}}_{i}}(\boldsymbol{\Omega}){\bar{\bar{\mathbf{U}}}_{i}}(\boldsymbol{\Omega})^{\prime}\right\}=\mathcal{B}(\boldsymbol{\Omega})$
Due to the consistency property of estimators for $\boldsymbol{\gamma}$, we have
$\sqrt{n}\left({\hat{{\boldsymbol{\Omega}}}}-\boldsymbol{\Omega}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}MVN\left\{\mathbf{0},\boldsymbol{\Sigma}(\boldsymbol{\Omega})\right\}$
where $\boldsymbol{\Sigma}(\boldsymbol{\Omega})=\mathcal{A}^{-1}(\boldsymbol{\Omega})\mathcal{B}(\boldsymbol{\Omega})\left\{\mathcal{A}^{-1}(\boldsymbol{\Omega})\right\}^{\prime}$ and
$\mathcal{A}(\boldsymbol{\Omega})=\left(\begin{matrix}\mathcal{A}_{11}&\mathcal{A}_{12}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{0}&\mathcal{A}_{22}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \mathcal{A}_{31}&\boldsymbol{0}&\mathcal{A}_{33}&\mathcal{A}_{34}&\mathcal{A}_{35}&\mathcal{A}_{36}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\mathcal{A}_{44}&\boldsymbol{0}&\boldsymbol{0}\\ \mathcal{A}_{51}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\mathcal{A}_{55}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\mathcal{A}_{66}\end{matrix}\right).$
D.4 Sandwich Variance Estimator and Wald Inference
For inference, we replace the population expectations in $\mathcal{A}(\cdot)$ and $\mathcal{B}(\cdot)$ with their empirical counterparts, yielding the sandwich covariance estimator
[…]
where
[…]
$W=\frac{\hat{\gamma}_{11}}{\text{s.e.}(\hat{\gamma}_{11})},\quad\text{s.e.}(\hat{\gamma}_{11})=\sqrt{\hat{\boldsymbol{\Sigma}}(\hat{\boldsymbol{\Omega}})_{[2,2]}},$
with an analogous procedure for $\gamma_{21}$ and other parameters of interest.
Appendix E. Proof of the Bias Induced by Misspecified Propensity Scores in Regression
Here, we aim to show the derivations to evaluate the form of the limiting values of estimators of causal effects under the two-stage approach, especially apply propensity score regression adjustment in both stages.
The response model is in the form of
$Y=\theta_{0}+{\theta}_{11}A+{\theta}_{12}A(D-c)+\mathbf{X}^{\prime}\boldsymbol{\theta}_{2}+E$
where $E\sim N(0,\tau^{2})$ and $(A,D,\mathbf{X})\perp E$.
E.1 Bias Caused by Misspecified Propensity Scores in Stage I
In the first stage, we let $\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})$ denote the Stage I propensity score under misspecification, and consider the regression adjustment
$E(Y|A=1,D,\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1});\tilde{\boldsymbol{\gamma}}_{1})=\tilde{\gamma}_{10}+\tilde{\gamma}_{11}D+\tilde{\gamma}_{12}\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})=\mu_{11}(\tilde{\boldsymbol{\phi}}_{1}),$
where $\tilde{\boldsymbol{\gamma}}_{1}=(\tilde{\gamma}_{10},\tilde{\gamma}_{11},\tilde{\gamma}_{12})^{\prime}$ and $\tilde{\boldsymbol{\phi}}_{1}=(\tilde{\boldsymbol{\gamma}}_{1}^{\prime},\tilde{\boldsymbol{\alpha}}_{1}^{\prime})^{\prime}$. Here $\tilde{\boldsymbol{\gamma}}_{1}$ is the solution to
$E\left\{\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})\right\}=\mathbf{0}$ (E.1)
where
$\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})=A\,\frac{\partial\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})}{\partial\tilde{\boldsymbol{\gamma}}_{1}}\{Y-\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})\}.$
By applying the law of iterated expectations with respect to $A$, we obtain
$E\bigl\{\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})\bigr\}=E\left[E\left[A\frac{\partial\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})}{\partial\tilde{\boldsymbol{\gamma}}_{1}}\{Y-\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})\}\middle|A\right]\right].$
Since $A$ is binary,
$E\left[A\frac{\partial\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})}{\partial\tilde{\boldsymbol{\gamma}}_{1}}\{Y-\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})\}\,\middle|\,A\right]=AE\left[\frac{\partial\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})}{\partial\tilde{\boldsymbol{\gamma}}_{1}}\{Y-\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})\}\,\middle|A\right].$
Therefore,
$E\bigl\{\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})\bigr\}=P(A=1)\,E\left[\frac{\partial\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})}{\partial\tilde{\boldsymbol{\gamma}}_{1}}\{Y-\mu_{11}(\tilde{\boldsymbol{\phi}}_{1})\}\,\middle|\,A=1\right].$
Since $P(A=1)>0$, solving the estimating equation (E.1) is equivalent to solving
$E\left\{\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})\middle|A=1\right\}=0.$
Here, we let $Y^{*}$ represent $Y|A=1$, $D^{*}$ represent $D|A=1$, $\mathbf{X}^{*}$ represent $\mathbf{X}|A=1$, and $\mathcal{D}^{*}=(Y^{*},S^{*},\boldsymbol{X}^{*})$. Now, we aim to evaluate this expectation. Firstly, we take the expectation with respect to $Y^{*}$ given $D^{*}$ and $\mathbf{X}^{*}$, then we have
$E_{Y^{*}}\bigl\{\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})|D^{*},\mathbf{X}^{*}\bigr\}$ $=E_{Y^{*}|D^{*},\mathbf{X}^{*}}\bigl\{\mathcal{U}_{11}(\mathcal{D};\tilde{\boldsymbol{\phi}}_{1})\bigr\}$
$=\bar{\mathbf{V}}_{1}^{*}\bigl\{E(Y^{*}|D^{*},\mathbf{X}^{*};\boldsymbol{\theta})-\bar{\mathbf{V}}_{1}^{*^{\prime}}\boldsymbol{\gamma}_{1}\bigr\}$ (E.2)
where $\bar{\mathbf{V}}_{1}^{*}=\{1,D^{*},\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\}^{\prime}$.
Then we take the expectation with respect to $D^{*}|\mathbf{X}^{*}$. This gives a system of equations as follows:
$\theta_{0}^{*}+\theta_{12}E(D^{*}|\mathbf{X}^{*})+\boldsymbol{\theta}_{2}^{\prime}\mathbf{X}^{*}-\bigl\{\gamma_{10}+\gamma_{11}E(D^{*}|\mathbf{X}^{*})+\gamma_{12}\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}=0$ (E.3)
$\theta_{0}^{*}E(D^{*}|\mathbf{X}^{*})+\theta_{12}E(D^{*2}|\mathbf{X}^{*})+\boldsymbol{\theta}_{2}^{\prime}E\bigl\{(D^{*}|\mathbf{X}^{*})\mathbf{X}^{*}\bigr\}$-
$\left[\gamma_{10}E(D^{*}|\mathbf{X}^{*})+\gamma_{11}E(D^{*2}|\mathbf{X}^{*})+\gamma_{12}E\bigl\{\log(D^{*}|\mathbf{X}^{*})\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]=0$ (E.4)
$\theta_{0}^{*}\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})+\theta_{12}E(D^{*}|\mathbf{X}^{*})\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})+\boldsymbol{\theta}_{2}^{\prime}\mathbf{X}^{*}\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})$-
$\bigl\{\gamma_{10}\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})+\gamma_{11}E(D^{*}|\mathbf{X}^{*})\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})+\gamma_{12}\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})^{2}\bigr\}=0$ (E.5)
where $\theta_{0}^{*}=\theta_{0}+\theta_{11}-\theta_{12}c$. Next, we take expectation with respect to $\mathbf{X}^{*}$. From (E.3), we get
$\theta_{0}^{*}+\theta_{12}E(D^{*})+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X}^{*})-\left[\gamma_{10}+\gamma_{11}E(D^{*})+\gamma_{12}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]=0$ (E.6)
From (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
$\theta_{0}^{*}E(D^{*})+\theta_{12}E(D^{*2})+\boldsymbol{\theta}_{2}^{\prime}\{\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})+E(D^{*})E(\mathbf{X}^{*})\}$-
$\biggl[\gamma_{10}E(D^{*})+\gamma_{11}E(D^{*2})+\gamma_{12}\left[\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+E(D^{*})E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]\biggr]=0$ (E.7)
where $\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})=\bigl\{\zeta_{11}(\mathbf{X}^{*}),...,\zeta_{1k}(\mathbf{X}^{*})\bigr\}^{\prime}$ where $\zeta_{1j}(\mathbf{X}^{*})=\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}_{j}),\mathbf{X}^{*}\bigr\}$ for j=1,2,,kj=1,2,...,k.
From (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
$\theta_{0}^{*}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+\theta_{12}\left[\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*})\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+E(D^{*})E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]$
$+\boldsymbol{\theta}_{2}^{\prime}\bigl[\boldsymbol{\phi}_{1}(\mathbf{X}^{*})+E\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\}E(\mathbf{X}^{*})\bigr]$
$-\biggl[\gamma_{10}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+\gamma_{11}\left[\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*})\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+E(D^{*})E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]$
$+\gamma_{12}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})^{2}\bigr\}\biggr]=0$ (E.8)
where $\boldsymbol{\phi}_{1}(\mathbf{X}^{*})=\bigl\{\phi_{11}(\mathbf{X}^{*}),...,\phi_{1k}(\mathbf{X}^{*})\bigr\}^{\prime}$ where $\phi_{1j}(\mathbf{X}^{*})=\text{cov}\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1}),\mathbf{X}_{j}^{*}\bigr\}$ for j=1,2,,kj=1,2,...,k. Then from (E.6), we can obtain the form of $\gamma_{10}$:
$\gamma_{10}=\theta_{0}^{*}+\theta_{12}E(D^{*})+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X}^{*})-\gamma_{11}E(D^{*})-\gamma_{12}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}$
Then, we replace $\gamma_{10}$ in (E.1 Bias Caused by Misspecified Propensity Scores in Stage I)
$\theta_{0}^{*}E(D^{*})+\theta_{12}E(D^{*2})+\boldsymbol{\theta}_{2}^{\prime}\bigl\{\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})+E(D^{*})E(\mathbf{X}^{*})\bigr\}$-
$\biggl[\left[\theta_{0}^{*}+\theta_{12}E(D^{*})+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X}^{*})-\gamma_{11}E(D^{*})-\gamma_{12}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]E(D^{*})$
$+\gamma_{11}E(D^{*2})+\gamma_{12}\left[\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+E(D^{*})E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]\biggr]=0$
$\theta_{0}^{*}E(D^{*})+\theta_{12}E(D^{*2})+\boldsymbol{\theta}_{2}^{\prime}\bigl\{\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})+E(D^{*})E(\mathbf{X}^{*})\bigr\}$-
$\biggl[\left[\theta_{0}^{*}+\theta_{12}E(D^{*})+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X}^{*})-\gamma_{11}E(D^{*})-\gamma_{12}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]E(D^{*})$
$+\gamma_{11}E(D^{*2})+\gamma_{12}\left[\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+E(D^{*})E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]\biggr]=0$
After some algebra, we get
$\theta_{12}[E(D^{*2})-E(D^{*})^{2}]+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})-\gamma_{11}[E(D^{*2})-E(D^{*})^{2}]$
$-\gamma_{12}\text{cov}\bigl[E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr]=0.$ (E.9)
After replacing $\gamma_{10}$ in (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
$\theta_{0}^{*}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+\theta_{12}\left[\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+E(D^{*})E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]$
$+\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})+E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}E(\mathbf{X}^{*})\right]$
$-\biggl[\left[\theta_{0}^{*}+\theta_{12}E(D^{*})+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X}^{*})-\gamma_{11}E(D^{*})-\gamma_{12}E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}$
$+\gamma_{11}\left[\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+E(D^{*})E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\right]+\gamma_{12}E\bigl[\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})^{2}\bigr]\biggr]=0.$ (E.10)
Similarly, after some algebra, from (E.1 Bias Caused by Misspecified Propensity Scores in Stage I) we finally get
$\theta_{12}\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})-\gamma_{11}\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}$
$-\gamma_{12}\left\{E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})^{2}\bigr\}-E\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}^{2}\right\}=0$
Then we can obtain $\gamma_{12}$ expressed as
$\gamma_{12}=\frac{\left[(\theta_{12}-\gamma_{11})\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})\right]}{\rm var\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}}.$ (E.11)
By replacing $\gamma_{12}$ in (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
$\theta_{12}\{E(D^{*2})-E(D^{*})^{2}]+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})-\gamma_{11}\{E(D^{*2})-E(D^{*})^{2}\}$
$-\frac{\left[(\theta_{12}-\gamma_{11})\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})\right]}{\rm var\bigl\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}}\text{cov}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}=0.$
After simplifying, we obtain
$\gamma_{11}$ $=\theta_{12}+\frac{\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})-\boldsymbol{\zeta}_{1}(\mathbf{X}^{*})\text{corr}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}\sqrt{\frac{\rm var\{E(D^{*}|\mathbf{X}^{*})\}}{\rm var\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\}}}\right]}{\rm var(D^{*})-\rm var\bigl\{E(D^{*}|\mathbf{X}^{*})\bigr\}\text{corr}^{2}\bigl\{E(D^{*}|\mathbf{X}^{*}),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}}$
$=\theta_{12}+\frac{\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)-\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)\text{corr}\bigl[E(D|\mathbf{X},A=1),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr]\sqrt{\frac{\rm var\{E(D|\mathbf{X},A=1)\}}{\rm var\{\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\}}}\right]}{\rm var(D|A=1)-\rm var\bigl\{E(D|\mathbf{X},A=1)\bigr\}\text{corr}^{2}\bigl\{E(D|\mathbf{X},A=1),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}}$
To make the form of the limiting value more clear, we rewrite it in the form of
$\gamma_{11}=\theta_{12}+\frac{\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)-\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)\rho\sqrt{\frac{\rm var\{E(D|\mathbf{X},A=1)\}}{\rm var\{\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})\}}}\right]}{\rm var(D|A=1)-\rm var\bigl\{E(D|\mathbf{X},A=1)\bigr\}\rho^{2}}$
where $\rho=\text{corr}\bigl\{E(D|\mathbf{X},A=1),\tilde{S}_{1}(\mathbf{X}^{*};\tilde{\boldsymbol{\alpha}}_{1})\bigr\}$.
E.2 Bias Caused by Misspecified Propensity Scores in Stage II
Let $\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$ be the Stage II propensity score under misspecification. We fit the propensity score regression adjustment model
$E\left\{Y|D,A,\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2});\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\gamma}}_{2}\right\}=\tilde{\gamma}_{20}+\tilde{\gamma}_{21}A+\tilde{\gamma}_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})+\text{offset}\left\{\tilde{\gamma}_{11}A(D-c)\right\}$
where $\tilde{\boldsymbol{\gamma}}_{2}=(\tilde{\gamma}_{21},\tilde{\gamma}_{22},\tilde{\gamma}_{23})^{\prime}$. The estimating equation for $\tilde{\boldsymbol{\gamma}}_{2}$ is
$E\left\{\mathcal{U}_{21}(D;\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\phi}}_{2})\right\}=\mathbf{0}$ (E.12)
where
$\mathcal{U}_{21}(D;\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\phi}}_{2})=\frac{\partial\mu_{21}(\tilde{\boldsymbol{\phi}}_{2})}{\partial\tilde{\boldsymbol{\phi}}_{2}}\left[Y-\left\{\mu_{21}(\tilde{\boldsymbol{\phi}}_{2})+\tilde{\gamma}_{11}A(D-c)\right\}\right]$
with $\mu_{21}(\tilde{\boldsymbol{\phi}}_{2})=\tilde{\gamma}_{20}+\tilde{\gamma}_{21}A+\tilde{\gamma}_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$.
We let $\bar{\mathbf{V}}_{2}=\{1,A,\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\}^{\prime}$, the estimating function for $\boldsymbol{\gamma}_{2}$ is expressed as
$\mathcal{U}_{21}(D;\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\phi}}_{2})=\bar{\mathbf{V}}_{2}\bigl\{Y-\tilde{\gamma}_{11}A(D-c)-\bar{\mathbf{V}}_{2}^{\prime}\boldsymbol{\gamma}_{2}\bigr\}.$
We obtain $\hat{\boldsymbol{\gamma}}_{2}$ by solving the equation
$E\{\mathcal{U}_{21}(D;\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\phi}}_{2})\}=\mathbf{0}$
Firstly, we take expectation with respect to $Y|D,A,\mathbf{X}$ and we get
$E_{Y|D,A,\mathbf{X}}\bigl\{\mathcal{U}_{21}(D;\tilde{\boldsymbol{\gamma}}_{1},\tilde{\boldsymbol{\phi}}_{2})\bigr\}=\bar{\mathbf{V}}_{2}\bigl\{E(Y|D,A,\mathbf{X};\boldsymbol{\theta})-\tilde{\gamma}_{11}A(D-c)-\bar{\mathbf{V}}_{2}^{\prime}\boldsymbol{\gamma}_{2}\bigr\}.$ (E.13)
Then we take expectation with respect to $D,A|\mathbf{X}$ and we can get a $3\times 1$ vector expressed as
$\theta_{0}+\theta_{11}E(A|\mathbf{X})+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(AD|\mathbf{X})-cE(A|\mathbf{X})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\mathbf{X}$
$-\{\gamma_{20}+\gamma_{21}E(A|\mathbf{X})+\gamma_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\}=0$ (E.14)
$\theta_{0}E(A|\mathbf{X})+\theta_{11}E(A^{2}|\mathbf{X})+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(A^{2}D|\mathbf{X})-cE(A^{2}|\mathbf{X})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}E\bigl\{(A|\mathbf{X})\mathbf{X}\bigr\}$
$-\left[\gamma_{20}E(A|\mathbf{X})+\gamma_{21}E(A^{2}|\mathbf{X})+\gamma_{22}E\bigl\{(A|\mathbf{X})\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right]=0$ (E.15)
$\theta_{0}{S}_{1}(\mathbf{X})+\theta_{11}E\bigl\{(A|\mathbf{X})\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(AD|\mathbf{X})\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}-cE\bigl\{(A|\mathbf{X})\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$
$+\boldsymbol{\theta}_{2}^{\prime}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\mathbf{X}-\left[\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\gamma_{20}+\gamma_{21}E\bigl\{(A|\mathbf{X})\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+\gamma_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})^{2}\right]=0.$ (E.16)
Our next step is to take expectation with respect to $\mathbf{X}$. From (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) we get
$\theta_{0}+\theta_{11}E(A)+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(AD)-cE(A)\bigr\}+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X})-\bigl\{\gamma_{20}+\gamma_{21}E(A)+\gamma_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}=0$ (E.17)
From (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) we get
$\theta_{0}E(A)+\theta_{11}E(A^{2})+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(A^{2}D)-cE(A^{2})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\bigl\{\boldsymbol{\zeta}_{2}(\mathbf{X})+E(A)E(\mathbf{X})\bigr\}$
$-\left[\gamma_{20}E(A)+\gamma_{21}E(A^{2})+\gamma_{22}\bigl[\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(A)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\bigr]\right]=0$ (E.18)
where $\boldsymbol{\zeta}_{2}(\mathbf{X})=\bigl\{\zeta_{21}(\mathbf{X}),\ldots,\zeta_{2k}(\mathbf{X})\bigr\}^{\prime}$ where $\zeta_{2j}(\mathbf{X})=\text{cov}\bigl\{E(A|\mathbf{X}),\mathbf{X}_{j}\bigr\}$ for j=1,2,,kj=1,2,...,k.
From (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) we get
$\theta_{0}E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+\theta_{11}\left[\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(A)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right]$
$+(\theta_{12}-\tilde{\gamma}_{11})\biggl[\text{cov}\bigl\{AD|\mathbf{X},\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(AD)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$
$-c\Bigl[\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(A)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\Bigr]\biggr]$
$+\boldsymbol{\theta}_{2}^{\prime}\left[\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}\bigr\}+E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}E(\mathbf{X})\right]$
$-\left[\gamma_{20}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})+\gamma_{21}\bigl[\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(A)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\bigr]+\gamma_{22}E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})^{2}\bigr\}\right]=0.$ (E.19)
Then from (E.17), we obtain $\gamma_{20}$:
$\gamma_{20}=\theta_{0}+\theta_{11}E(A)+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(AD)-cE(A)\bigr\}+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X})-\gamma_{21}E(A)-\gamma_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})$
We replace $\gamma_{20}$ in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) and get:
$\theta_{0}E(A)+\theta_{11}E(A^{2})+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(A^{2}D)-cE(A^{2})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\bigl\{\boldsymbol{\zeta}_{2}(\mathbf{X})+E(A)E(\mathbf{X})\bigr\}$
$-\left[\theta_{0}+\theta_{11}E(A)+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(AD)-cE(A)\bigr\}+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X})-\gamma_{21}E(A)-\gamma_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\right]E(A)$
$-\gamma_{21}E(A^{2})-\gamma_{22}\left[\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(\mathbf{X})E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right]=0.$
After algebra, we get
$\theta_{11}\bigl\{E(A^{2})-E(A)^{2}\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{2}(\mathbf{X})-\gamma_{21}\bigl\{E(A^{2})-E(A)^{2}\bigr\}-\gamma_{22}\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$
$+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(A^{2}D)-E(AD)E(A)+cE(A)^{2}-cE(A^{2})\bigr\}=0.$ (E.20)
To solve for $\gamma_{22}$, we replace $\gamma_{20}$ in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) to get
$\theta_{0}E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+\theta_{11}\left[\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(A)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right]$
$+(\theta_{12}-\tilde{\gamma}_{11})\Bigl[\text{cov}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(AD)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$
$-c\bigl[\text{cov}\big\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+E(A)E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\bigr]\Bigr]$
$+\boldsymbol{\theta}_{2}^{\prime}\left[\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}\bigr\}+E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}E(\mathbf{X})\right]$
$-\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\left[\theta_{0}+\theta_{11}E(A)+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(AD)-cE(A)\bigr\}+\boldsymbol{\theta}_{2}^{\prime}E(\mathbf{X})-\gamma_{21}E(A)-\gamma_{22}\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\right]$
$-\gamma_{21}\left[\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right]-\gamma_{22}E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})^{2}\bigr\}=0.$
We simplify it to get
$\theta_{11}\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\text{cov}\bigl\{E(\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})|\mathbf{X}),\mathbf{X}\bigr\}-\gamma_{21}\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$
$-\gamma_{22}\left[E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})^{2}\bigr\}-E\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}^{2}\right]$
$+(\theta_{12}-\tilde{\gamma}_{11})\left[\text{cov}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}-c\cdot\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right]=0.$
Then we can obtain $\gamma_{22}$ expressed by
$\gamma_{22}=\frac{(\theta_{11}-\gamma_{21})\delta_{1}+\boldsymbol{\theta}_{2}^{\prime}\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}\bigr\}+(\theta_{12}-\tilde{\gamma}_{11})(\delta_{2}-c\delta_{1})}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}$ (E.21)
where $\delta_{1}=\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$ and $\delta_{2}=\text{cov}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$. Next, we substitute $\gamma_{22}$ in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II):
$\theta_{11}\bigl\{E(A^{2})-E(A)^{2}\bigr\}+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{2}(\mathbf{X})-\gamma_{21}\bigl\{E(A^{2})-E(A)^{2}\bigr\}$
$-\frac{(\theta_{11}-\gamma_{21})\delta_{1}+\boldsymbol{\theta}_{2}^{\prime}\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}\bigr\}+(\theta_{12}-\tilde{\gamma}_{11})(\delta_{2}-c\delta_{1})}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\delta_{1}$
$+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{E(A^{2}D)-E(AD)E(A)+cE(A)^{2}-cE(A^{2})\bigr\}=0.$
By simplifying, we get
$\theta_{11}\rm var(A)+\boldsymbol{\theta}_{2}^{\prime}\boldsymbol{\zeta}_{2}(\mathbf{X})-\gamma_{21}\rm var(A)$
$-\frac{(\theta_{11}-\gamma_{21})\delta_{1}+\boldsymbol{\theta}_{2}^{\prime}\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}\bigr\}+(\theta_{12}-\tilde{\gamma}_{11})(\delta_{2}-c\delta_{1})}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\delta_{1}$
$+(\theta_{12}-\tilde{\gamma}_{11})\bigl\{\text{cov}(AD,A)-c\rm var(A)\bigr\}=0.$
Then we re-organize it and rewrite it as
$\theta_{11}\left[\rm var(A)-\frac{\delta_{1}^{2}}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\right]+\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{2}(\mathbf{X})-\frac{\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}\bigr\}\delta_{1}}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\right]$
$+(\theta_{12}-\tilde{\gamma}_{11})\left[\text{cov}(AD,A)-c\rm var(A)-\frac{\delta_{1}\delta_{2}-c\delta_{1}^{2}}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\right]=\gamma_{21}\left[\rm var(A)-\frac{\delta_{1}^{2}}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\right].$ (E.22)
Here we simplify some terms in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II),
$\theta_{11}\left[\rm var(A)-\frac{\delta_{1}^{2}}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\right]=\theta_{11}\left[\rm var(A)-\rm var\bigl\{E(A|\mathbf{X})\bigr\}\text{corr}^{2}\bigl\{E(A|\mathbf{X})\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right],$
$(\theta_{12}-\tilde{\gamma}_{11})\left[\text{cov}(AD,A)-c\rm var(A)-\frac{\delta_{1}\delta_{2}-c\delta_{1}^{2}}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\right]$
$=(\theta_{12}-\tilde{\gamma}_{11})\biggl[\text{cov}(AD,A)-cE\bigl\{\rm var(A|\mathbf{X})\bigr\}\biggr$.
$\left.-\delta_{1}\Bigl[\sqrt{\rm var\bigl\{E(A|\mathbf{X})\bigr\}\rm var\{E(AD|\mathbf{X})\}}\text{corr}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\Bigr]\right]$
$\gamma_{21}\left[\rm var(A)-\frac{\delta_{1}^{2}}{\rm var\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}}\right]=\gamma_{21}\left[\rm var(A)-\rm var\bigl\{E(A|\mathbf{X})\bigr\}\text{corr}^{2}\bigl\{E(A|\mathbf{X})\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}\right].$
Here we let
$\rho_{21}=\text{corr}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\};\rho_{22}=\text{corr}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\};$
$\boldsymbol{\zeta}_{2}(\mathbf{X})=\bigl\{\zeta_{21}(\mathbf{X}),...,\zeta_{2k}(\mathbf{X})\bigr\}^{\prime}$
where $\zeta_{2j}(\mathbf{X})=\text{cov}\bigl\{E(A|\mathbf{X}),\mathbf{X}_{j}\bigr\}.$ Let
$\boldsymbol{\phi}_{2}(\mathbf{X})=\bigl\{\phi_{21}(\mathbf{X}),...,\phi_{2k}(\mathbf{X})\bigr\}^{\prime}$
where $\phi_{2j}(\mathbf{X})=\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}_{j}\bigr\}$ for j=1,2,,kj=1,2,...,k. Then (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) can be expressed as
$\gamma_{21}=\theta_{11}+\frac{\boldsymbol{\theta}_{2}^{\prime}\Bigl[\boldsymbol{\zeta}_{2}(\mathbf{X})-\frac{\boldsymbol{\phi}_{2}(\mathbf{X})\delta_{1}}{\rm var\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\}}\Bigr]}{\rm var(A)-\rm var\{E(A|\mathbf{X})\}\rho_{21}^{2}}$
$+\frac{(\theta_{12}-\tilde{\gamma}_{11})\biggl[\text{cov}(AD,A)-cE\bigl\{\rm var(A|\mathbf{X})\bigr\}-\delta_{1}\Bigl[\sqrt{\rm var\{E(A|\mathbf{X})\}\rm var\{E(AD|\mathbf{X})\}}\rho_{22}\Bigr]\biggr]}{\rm var(A)-\rm var\bigl\{E(A|\mathbf{X})\bigr\}\rho_{21}^{2}}$
and we can obtain $\gamma_{21}$ expressed as
$\gamma_{21}=\theta_{11}+\frac{\boldsymbol{\theta}_{2}^{\prime}\left[\boldsymbol{\zeta}_{2}(\mathbf{X})-\boldsymbol{\phi}_{2}(\mathbf{X})\rho_{21}\sqrt{\frac{\rm var\{E(A|\mathbf{X})\}}{\rm var\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\}}}\right]}{\rm var(A)-\rm var\bigl\{E(A|\mathbf{X})\bigr\}\rho_{21}^{2}}$
$+\frac{(\theta_{12}-\tilde{\gamma}_{11})\left[\text{cov}(AD,A)-cE\bigl\{\rm var(A|\mathbf{X})\bigr\}-\delta_{1}\rho_{22}\sqrt{\rm var\{E(A|\mathbf{X})\}\rm var\{E(AD|\mathbf{X})\}}\right]}{\rm var(A)-\rm var\{E(A|\mathbf{X})\}\rho_{21}^{2}}.$
Appendix F. Additional Materials on Simulation Study
F.1 Model Misspecification
To explore the performance of estimators under misspecification we consider the following types of misspecified models:
  • $S_{1}$: Misspecified $S_{1}$ model–omit $X_{12}$ from (2) and omit $X_{2}$ from (26)
  • $S_{2}$: Misspecified $S_{2}$ model–omit $X_{22}$ from (3) and omit $X_{2}$ from (27)
  • $m_{a}(\mathbf{X};\boldsymbol{\theta})$: Misspecified imputation model–omit $X_{12}$ from (1), and omit $X_{2}$ from (28)
  • $S_{1}+m_{a}(\mathbf{X};\boldsymbol{\theta})$: Misspecified $S_{1}$ model and imputation model
  • $S_{2}+m_{a}(\mathbf{X};\boldsymbol{\theta})$: Misspecified $S_{2}$ model and imputation model
F.2 Additional Simulation Results
Table Table 1: Empirical results for estimators of causal effects with $\pi^{*}=0.25$ under six methods of analysis based on correctly specified models.
Minimal PS: $\mathbf{S}=\{S_{1}(\mathbf{Z}_{1}),S_{2}(\mathbf{Z}_{2})\}^{\prime}$} Expanded PS: $\mathbf{S}=\{S_{1}(\mathbf{X}),S_{2}(\mathbf{X})\}^{\prime}$}
Method Exposure Effect Ebias ESE RSE ECP(%) Ebias ESE RSE ECP(%)
Data Generation Model 1
Naive $D$ $\theta_{12}$ 0.762 0.050 0.049 0.0 - - - -
AA $\theta_{11}$ 0.971 0.134 0.134 0.0 - - - -
DG model $D$ $\theta_{12}$ -0.001 0.034 0.033 94.4 - - - -
AA $\theta_{11}$ -0.004 0.076 0.077 94.8 - - - -
Two PS Reg $D$ $\theta_{12}$ -0.023 0.042 0.042 92.0 -0.023 0.041 0.041 91.7
AA $\theta_{11}$ -0.008 0.084 0.085 94.8 -0.007 0.078 0.078 94.7
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ -0.002 0.085 0.083 94.1 -0.003 0.064 0.063 94.4
AA $\theta_{11}$ -0.004 0.112 0.112 95.1 -0.005 0.080 0.081 95.2
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ -0.002 0.085 0.083 94.1 -0.003 0.064 0.063 94.4
AA $\theta_{11}$ -0.004 0.160 0.152 94.0 -0.006 0.138 0.130 93.1
$\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$ $D$ $\theta_{12}$ -0.002 0.085 0.083 94.1 -0.003 0.064 0.063 94.4
AA $\theta_{11}$ -0.004 0.092 0.092 94.9 -0.005 0.089 0.089 95.0
Data Generation Model 2
Naive $D$ $\theta_{12}$ 0.463 0.049 0.048 0.0 - - - -
AA $\theta_{11}$ 0.496 0.090 0.089 0.0 - - - -
DG model $D$ $\theta_{12}$ 0.001 0.045 0.044 93.7 - - - -
AA $\theta_{11}$ 0.001 0.079 0.077 94.3 - - - -
Two PS Reg $D$ $\theta_{12}$ 0.001 0.045 0.049 96.9 - - - -
AA $\theta_{11}$ 0.001 0.080 0.080 94.8 - - - -
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.001 0.065 0.063 93.8 - - - -
AA $\theta_{11}$ 0.002 0.081 0.078 94.3 - - - -
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.001 0.065 0.063 93.8 - - - -
AA $\theta_{11}$ 0.003 0.101 0.096 93.7 - - - -
$\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.001 0.065 0.063 93.8 - - - -
AA $\theta_{11}$ 0.002 0.091 0.088 94.5 - - - -
  • DG model: Data generation model; Two PS Reg: Two-dimensional propensity score regression adjustment; $\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$: Using PS regression adjustment in Stage I and II; $\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$: Using PS regression adjustment in Stage I and IPW in Stage II; $\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$: Using PS regression adjustment in Stage I and AIPW in Stage II.
Table Table 2: Empirical results for estimators of causal effects with $\pi^{*}=0.25$ for two-propensity-score regression adjustment and two-stage analysis under model misspecification.
Misspecified $S_{1}$ Misspecified $S_{2}$
Method Exposure Effect Ebias ESE RSE ECP(%) Ebias ESE RSE ECP(%)
Data Generation Model 1
Two PS Reg $D$ $\theta_{12}$ 0.082 0.047 0.051 64.2 -0.014 0.042 0.047 96.2
AA $\theta_{11}$ 0.008 0.097 0.098 95.0 0.243 0.086 0.087 18.4
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.349 0.090 0.088 2.7 -0.002 0.085 0.083 94.1
AA $\theta_{11}$ 0.050 0.105 0.106 92.4 0.335 0.113 0.114 15.7
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.349 0.090 0.090 2.7 -0.002 0.085 0.083 94.1
AA $\theta_{11}$ 0.135 0.140 0.134 77.3 0.338 0.126 0.126 24.2
Data Generation Model 2
Two PS Reg $D$ $\theta_{12}$ -0.098 0.065 0.062 65.3 0.001 0.065 0.063 93.7
AA $\theta_{11}$ -0.104 0.087 0.084 75.2 0.192 0.090 0.087 39.8
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.145 0.061 0.060 30.8 0.001 0.065 0.063 93.8
AA $\theta_{11}$ 0.027 0.081 0.078 92.7 0.247 0.081 0.079 13.1
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.145 0.061 0.060 30.8 0.001 0.065 0.063 93.8
AA $\theta_{11}$ 0.070 0.095 0.090 85.4 0.248 0.084 0.082 16.3
  • $S_{1}$ represents the propensity score model for the continuous exposure $D$ conditioned on $A=1$, given by $S_{1}=E(D|A=1,\mathbf{Z}_{1})$; $S_{2}$ represents the propensity score model for the binary exposure $A$, given by $S_{2}=E(A|\mathbf{Z}_{2})$.
Table Table 3: Empirical results for estimators of effects for both parts with $\pi^{*}=0.25$ for applying PS regression adjustment in Stage I and AIPW in Stage II under model misspecification.
Misspecified Model Exposure Effect Ebias ESE RSE ECP(%)
Data Generation Model 1
$S_{1}$ $D$ $\theta_{12}$ 0.349 0.090 0.088 2.7
AA $\theta_{11}$ 0.135 0.101 0.100 73.3
$S_{2}$ $D$ $\theta_{12}$ -0.002 0.085 0.083 94.1
AA $\theta_{11}$ -0.011 0.090 0.090 94.8
$m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ -0.002 0.085 0.083 94.1
AA $\theta_{11}$ -0.001 0.096 0.100 95.9
$S_{1}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.349 0.090 0.088 2.7
AA $\theta_{11}$ 0.138 0.104 0.107 74.4
$S_{2}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ -0.002 0.085 0.083 94.1
AA $\theta_{11}$ 0.226 0.086 0.088 26.0
Data Generation Model 2
$S_{1}$ $D$ $\theta_{12}$ 0.145 0.061 0.060 30.8
AA $\theta_{11}$ 0.068 0.088 0.085 86.3
$S_{2}$ $D$ $\theta_{12}$ 0.001 0.065 0.063 93.8
AA $\theta_{11}$ 0.002 0.089 0.086 94.3
$m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.001 0.065 0.063 93.8
AA $\theta_{11}$ 0.003 0.094 0.096 95.5
$S_{1}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.145 0.061 0.060 30.8
AA $\theta_{11}$ 0.069 0.090 0.090 87.0
$S_{2}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.001 0.065 0.063 93.8
AA $\theta_{11}$ 0.247 0.083 0.081 15.0
  • $S_{1}$ represents the propensity score model for the continuous exposure $D$ conditioned on $A=1$, given by $S_{1}=E(D|A=1,\mathbf{Z}_{1})$; $S_{2}$ represents the propensity score model for the binary exposure $A$, given by $S_{2}=E(A|\mathbf{Z}_{2})$; $m_{a}(\mathbf{X};\boldsymbol{\theta})$ represents the imputation models for $a=0,1$.
Table Table 4: Empirical results for estimators of causal effects with $\pi^{*}=0.75$ under six methods of analysis based on correctly specified models.
Minimal PS: $\mathbf{S}=\{S_{1}(\mathbf{Z}_{1}),S_{2}(\mathbf{Z}_{2})\}^{\prime}$} Expanded PS: $\mathbf{S}=\{S_{1}(\mathbf{X}),S_{2}(\mathbf{X})\}^{\prime}$}
Method Exposure Effect Ebias ESE RSE ECP(%) Ebias ESE RSE ECP(%)
Data Generation Model 1
Naive $D$ $\theta_{12}$ 0.765 0.029 0.029 0.0 - - - -
AA $\theta_{11}$ 0.976 0.160 0.157 0.0 - - - -
DG model $D$ $\theta_{12}$ -0.001 0.026 0.027 95.9 - - - -
AA $\theta_{11}$ 0.002 0.077 0.078 94.9 - - - -
Two PS Reg $D$ $\theta_{12}$ 0.014 0.033 0.033 92.3 0.013 0.031 0.031 92.4
AA $\theta_{11}$ 0.007 0.084 0.085 95.1 0.007 0.078 0.079 94.8
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ -0.001 0.048 0.048 95.5 -0.001 0.036 0.036 95.6
AA $\theta_{11}$ 0.004 0.109 0.112 95.4 0.002 0.079 0.081 95.2
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ -0.001 0.048 0.048 95.5 -0.001 0.036 0.036 95.6
AA $\theta_{11}$ 0.007 0.150 0.147 93.7 0.004 0.132 0.126 94.3
$\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$ $D$ $\theta_{12}$ -0.001 0.048 0.048 95.5 -0.001 0.036 0.036 95.6
AA $\theta_{11}$ 0.001 0.083 0.085 95.3 0.001 0.084 0.085 95.1
Data Generation Model 2
Naive $D$ $\theta_{12}$ 0.465 0.028 0.028 0.0 - - - -
AA $\theta_{11}$ 0.498 0.099 0.097 0.15 - - - -
DG model $D$ $\theta_{12}$ 0.001 0.050 0.044 93.7 - - - -
AA $\theta_{11}$ 0.001 0.080 0.077 94.0 - - - -
Two PS Reg $D$ $\theta_{12}$ 0.0002 0.032 0.032 95.1 - - - -
AA $\theta_{11}$ 0.001 0.081 0.078 94.5 - - - -
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.001 0.037 0.037 95.0 - - - -
AA $\theta_{11}$ 0.002 0.082 0.079 94.1 - - - -
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.001 0.037 0.037 95.0 - - - -
AA $\theta_{11}$ 0.002 0.082 0.079 94.1 - - - -
$\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.001 0.037 0.037 95.0 - - - -
AA $\theta_{11}$ 0.001 0.087 0.084 94.4 - - - -
  • DG model: Data generation model; Two PS Reg: Two-dimensional propensity score regression adjustment; $\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$: Using PS regression adjustment in Stage I and II; $\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$: Using PS regression adjustment in Stage I and IPW in Stage II; $\text{PS}_{\text{I}}+\text{AIPW}_{\text{II}}$: Using PS regression adjustment in Stage I and AIPW in Stage II.
Table Table 5: Empirical results for estimators of causal effects with $\pi^{*}=0.75$ for two-propensity-score regression adjustment and two-stage analysis under model misspecification.
Misspecified $S_{1}$ Misspecified $S_{2}$
Method Exposure Effect Ebias ESE RSE ECP(%) Ebias ESE RSE ECP(%)
Data Generation Model 1
Two PS Reg $D$ $\theta_{12}$ 0.206 0.026 0.027 0.0 0.007 0.032 0.034 95.3
AA $\theta_{11}$ 0.077 0.070 0.070 80.7 0.255 0.085 0.087 15.2
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.350 0.050 0.051 0.0 -0.001 0.048 0.048 95.5
AA $\theta_{11}$ 0.134 0.109 0.112 78.5 0.344 0.120 0.114 13.2
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.350 0.050 0.051 0.0 -0.001 0.048 0.048 95.5
AA $\theta_{11}$ 0.053 0.150 0.157 90.7 0.347 0.118 0.123 19.6
Data Generation Model 2
Two PS Reg $D$ $\theta_{12}$ -0.011 0.036 0.036 93.4 0.001 0.037 0.037 95.1
AA $\theta_{11}$ 0.089 0.086 0.083 81.1 0.276 0.084 0.081 8.6
$\text{PS}_{\text{I}}+\text{PS}_{\text{II}}$ $D$ $\theta_{12}$ 0.146 0.036 0.035 1.6 0.001 0.037 0.036 95.0
AA $\theta_{11}$ 0.066 0.081 0.079 86.0 0.246 0.082 0.080 14.4
$\text{PS}_{\text{I}}+\text{IPW}_{\text{II}}$ $D$ $\theta_{12}$ 0.146 0.036 0.035 1.6 0.001 0.037 0.036 95.0
AA $\theta_{11}$ 0.024 0.099 0.093 92.0 0.247 0.084 0.081 16.0
  • $S_{1}$ represents the propensity score model for the continuous exposure $D$ conditioned on $A=1$, given by $S_{1}=E(D|A=1,\mathbf{Z}_{1})$; $S_{2}$ represents the propensity score model for the binary exposure $A$, given by $S_{2}=E(A|\mathbf{Z}_{2})$.
Table Table 6: Empirical results for estimators of effects for both parts with $\pi^{*}=0.75$ for applying PS regression adjustment in Stage I and AIPW in Stage II under model misspecification.
Misspecified Model Exposure Effect Ebias ESE RSE ECP(%)
Data Generation Model 1
$S_{1}$ $D$ $\theta_{12}$ 0.350 0.050 0.051 0.0
AA $\theta_{11}$ 0.048 0.087 0.089 92.0
$S_{2}$ $D$ $\theta_{12}$ -0.001 0.048 0.048 95.5
AA $\theta_{11}$ -0.020 0.087 0.089 95.1
$m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ -0.001 0.048 0.048 95.5
AA $\theta_{11}$ 0.005 0.087 0.088 95.0
$S_{1}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.350 0.050 0.051 0.0
AA $\theta_{11}$ 0.051 0.091 0.091 90.7
$S_{2}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ -0.001 0.048 0.048 95.5
AA $\theta_{11}$ 0.231 0.079 0.082 18.5
Data Generation Model 2
$S_{1}$ $D$ $\theta_{12}$ 0.146 0.036 0.034 1.6
AA $\theta_{11}$ 0.023 0.087 0.084 92.9
$S_{2}$ $D$ $\theta_{12}$ 0.001 0.037 0.036 95.0
AA $\theta_{11}$ 0.001 0.085 0.082 94.4
$m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.001 0.037 0.036 95.0
AA $\theta_{11}$ 0.001 0.090 0.086 93.5
$S_{1}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.146 0.036 0.035 1.6
AA $\theta_{11}$ 0.234 0.090 0.087 92.8
$S_{2}+m_{a}(\mathbf{X};\boldsymbol{\theta})$ $D$ $\theta_{12}$ 0.001 0.037 0.036 95.0
AA $\theta_{11}$ 0.246 0.083 0.080 15.1
  • $S_{1}$ represents the propensity score model for the binary exposure $A$, given by $S_{1}=E(A|\mathbf{Z}_{1})$; $S_{2}$ represents the propensity score model for the continuous exposure $D$ conditioned on $A=1$, given by $S_{2}=E(D|A=1,\mathbf{Z}_{2})$; $m_{a}(\mathbf{X};\boldsymbol{\theta})$ represents the imputation models for $a=0,1$.
We consider two scenarios where the probability of exposure is set to $\pi^{*}=P(A=1)=0.25$ and $0.75$. Table 1, Table 2, and Table 3 provide empirical results for estimators of causal effects under $\pi^{*}=0.25$. These results are consistent with those when $\pi^{*}=0.5$. The empirical biases for the effect in Stage II are large when either $S_{1}$ or $S_{2}$ is misspecified, leading to reduced performance. The double robustness property of the AIPW approach ensures consistency when only one of the models in Stage II is correctly specified. Table 4, Table 5, and Table 6 display the simulation results for $\pi^{*}=0.75$. The findings again reflect similar patterns observed under $\pi^{*}=0.5$ and $\pi^{*}=0.25$. When $S_{1}$ is misspecified in the first stage, the estimator performance suffers in both stages, indicating a significant sensitivity to model misspecification in the continuous part of the exposure.
Appendix G. Additional Information about the Application
G.1 Effects of Prenatal Covariates on PAE
Table Table 7: Estimated effects of prenatal covariates on the continuous and binary exposure components
$E(D|A=1,\mathbf{X})$ $P(A=1|\mathbf{X})$
Estimate Standard error Estimate Standard error
Intercept -4.183 1.003 -0.965 1.926
Smoking during pregnancy (cigarettes/day) 0.016 0.007 0.088 0.025
Socioeconomic status at birth -0.007 0.003 -0.003 0.005
Child’s gender (female) -0.165 0.165 -0.608 0.314
Mother’s age at delivery 0.057 0.018 0.056 0.036
Prenatal Beck score 0.026 0.011 0.026 0.023
HOME score at infancy -0.004 0.019 0.039 0.036
Biological mother’s PPVT -0.010 0.007 0.001 0.014
Biological mother’s marital status -0.042 0.276 -0.103 0.487
Prenatal cocaine exposure (days/month) -0.061 0.038 -0.102 0.070
Prenatal marijuana exposure (days/month) -0.012 0.032 0.220 0.198
Prenatal opiate exposure (days/month) 0.193 0.079 0.057 0.253
Number of prenatal visits -0.121 0.032 -0.021 0.061
Biological mother’s education (years) 0.098 0.059 -0.041 0.115
Parity -0.003 0.095 -0.471 0.220
Gravidity 0.036 0.057 0.271 0.145
Gestational age at screening 0.018 0.013 -0.010 0.025
G.2 Details of Multiple Imputation Procedure
In our data application, the Detroit dataset contains missing values in the outcome and covariates. 181 women have at least one missing value (41 with missing WISC-III and 159 with at least one missing covariate, with overlap between these groups). We address this by generating 20 multiply imputed datasets using an imputation model including all variables;Multiple imputation relies on the assumption that data are missing at random (MAR), which means the probability of missingness may depend on observed data but not on the unobserved values themselves. Following this assumption, we use the mice package in R to perform multiple imputation by chained equations (MICE) and generate $20$ imputed datasets [Van Buuren and Groothuis-Oudshoorn, 2011]. The imputation model included all variables used in the outcome model, the two propensity score models, and several auxiliary variables associated with the outcome or with missingness, following the recommendations in White et al. [2011] and Van Buuren and Van Buuren [2012].
Specifically, the imputation model included the outcome variable: the WISC-III Freedom from Distractibility Index measured at child age 7; the exposure variables: drinking status and average daily absolute alcohol consumption (AA/day) during pregnancy; and all covariates used in the propensity score models. In addition, auxiliary variables, which are the proportion of drinking days and the amount of alcohol consumed per drinking occasion across pregnancy are included. As shown in Li et al. [2023], these measures of prenatal alcohol exposure reflect distinct patterns of drinking behaviour and have differing impacts on children’s cognition; they also exhibit different relationships with covariates when constructing propensity score models. Therefore, including them as auxiliary variables enhances the plausibility of the MAR assumption and improves the accuracy of imputed values [White et al., 2011].
Two binary variables, child’s gender and mother’s marital status, are imputed using logistic regression. All other variables are continuous and imputed using predictive mean matching (PMM). The default chained equations algorithm in mice is run for 5 iterations per imputation.
After imputation, the analysis under the proposed methods is described in Section 6 of the main text. The resulting point estimates and standard errors are then combined using Rubin’s Rules to obtain valid overall estimates and associated uncertainty [Rubin, 1976].
G.3 Details of Method Implementation in Data Application
G.3.1 Conventional Covariate Regression Adjustment.
Let $A$ be a binary exposure representing whether the mother drinks during pregnancy, $T$ denote the absolute alcohol consumption per day (in ounces), and $D=\log T$ represent the logarithm of daily alcohol consumption if $A=1$. The vector of all confounders is denoted by $\mathbf{X}$. The regression model we consider is:
$E(Y_{i}\mid A_{i}=1,D_{i},\mathbf{X}_{i})=\theta_{0}+\theta_{11}A_{i}+\theta_{12}A_{i}(D_{i}-c)+\mathbf{X}^{\prime}_{i}\boldsymbol{\theta}_{2},$
where $c$ is a fixed constant representing the mean logarithm of daily alcohol consumption (in ounces). In this analysis, we set c=2.31c=-2.31 which corresponds to the empirical mean value of $D$ among mothers in this study who reported drinking.
G.3.2 Two-Propensity-Score Regression Adjustment.
Firstly, we define two propensity score models:
$S_{i1}=E(D|A=1,\boldsymbol{X}_{i})=\alpha_{20}+\boldsymbol{\alpha}_{21}\boldsymbol{X}_{i},$
$S_{i2}=E(A|\boldsymbol{X}_{i})=\text{expit}(\alpha_{10}+\boldsymbol{\alpha}_{11}\boldsymbol{X}_{i}).$
Then the two–propensity-score regression adjustment is specified by the linear model
$E(Y_{i}|A_{i}=1,D_{i},\mathbf{S}_{i})=\eta_{0}+{\eta}_{11}A_{i}+{\eta}_{12}A_{i}(D_{i}-c)+{\eta}_{21}S_{i1}+\eta_{22}S_{i2}$
where $\mathbf{S}_{i}=(S_{i1},S_{i2})^{\prime}$.
G.3.3 Two-stage Approach.
We assess the causal effects with a semi-continuous exposure by a two-stage approach. In the first stage, we aim to obtain the effect of the logarithm of the absolute alcohol consumption by propensity score regression adjustment. Then we treat $\hat{\gamma}_{11}A(D-c)$ as an offset term where $\hat{\gamma}_{11}$ is the estimated causal effect of alcohol consumption. In stage II, we apply propensity score regression adjustment, IPW, and AIPW approaches to find the effect of prenatal drinking status.